### v1

# Optimization under Attack: Resilience, Vulnerability, and the Path to Collapse

Abstract
Optimization is instrumental for improving operations of large-scale socio-technical infrastructures of Smart Cities, for instance, energy and traffic systems. In particular, understanding the performance of multi-agent discrete-choice combinatorial optimization under distributed adversary attacks is a compelling and underexplored problem, since multi-agent systems exhibit a large number of remote control variables that can influence in an unprecedented way the cost-effectiveness of distributed optimization heuristics. This paper unravels for the first time the trajectories of distributed optimization from resilience to vulnerability, and finally to collapse under varying adversary influence. Using real-world data to emulate over 28 billion multi-agent optimization scenarios, we exhaustively assess how the number of agents with different adversarial severity and network positioning influences optimization performance, including the influence on Pareto optimal points. With this novel large-scale dataset, made openly available as a benchmark, we disentangle how optimization remains resilient to adversaries and which adversary conditions are required to make optimization vulnerable or collapsed. These new findings can provide new insights for designing self-healing strategies for fault-tolerance and fault-correction in adversarial distributed optimization that have been missing so far.
Index Terms:
Optimization, Multi-agent systems, Adversary behavior, Resilience, Vulnerability, Distributed systems, Fault-tolerance
I Introduction
The rapid development of Internet of Things (IoT) applications has brought about transformative changes in numerous domains, ranging from smart cities to industrial automation and healthcare [1]. These applications involve vast networks of interconnected devices that generate substantial amounts of data, and require efficient and decentralized decision-making for processing [2]. Distributed optimization is essential in such scenarios, as it enables multiple agents to collaborate effectively without centralized coordination. This approach ensures scalability, reliability, and robust performance across diverse applications, including energy management, autonomous systems, and federated learning [3, 4, 5, 6, 7].
While distributed optimization is essential for achieving system-wide objectives, most algorithms rely on assumptions of rational, cooperative, and non-adversarial agents contributing toward the global objective without prioritizing their individual goals over collective outcomes. However, real-world scenarios often deviate from these assumptions, as adversarial agents can disrupt optimization systems by introducing inaccuracies, manipulating decisions, or compromising functionality [8, 9, 10, 11]. For example, in smart grid systems, where autonomous agents (e.g., households) collaborate to manage energy distribution and prevent power outages, certain households may act adversarially by manipulating consumption data and prioritizing self-interest, i.e. their thermal comfort. Such disruptions distort energy allocation processes, leading to inefficient resource use and blackouts [9, 10, 12, 13, 14].
Several studies have examined multi-agent optimization in continuous-choice frameworks [8, 9, 10, 11, 13, 15]. While stochastic optimization techniques address noise, they lack mechanisms to counteract strategic manipulations by adversarial agents [16, 17]. These manipulations amplify system vulnerability by prioritizing individual goals over collective objectives, which can lead to inefficiencies, instability, and eventual collapse [18, 19]. Despite this, adversarial disruptions in discrete-choice settings have received little attention.
This paper addresses this gap by investigating the resilience, vulnerability, and collapse of multi-agent systems in discrete-choice optimization under adversarial conditions, offering a novel framework for understanding system behavior. Resilience refers to the system ability to maintain performance despite adversarial influence. Vulnerability captures the emergence of inefficiencies due to adversarial behaviors, and collapse occurs when the system fails to achieve its objectives. This study evaluates the impact of adversarial agents on optimization performance, analyzing critical parameters such as the number of adversarial agents, their adversarial severity, and their network positions. By systematically examining these dynamics, the paper provides critical insights into the conditions under which systems transition from resilience to collapse, unravelling new insights for the development of strategies [20] for self-healing, fault-tolerance, and fault-correction in adversarial environments.
The main contributions of this paper are outlined as follows:
  • An adversarial model for discrete-choice multi-objective optimization problems.
  • A novel evaluation framework for characterizing resilience, vulnerability, and collapse in distributed optimization systems under adversarial conditions.
  • The applicability of the proposed adversarial model in different application domains.
  • Open large-scale benchmark datasets comprising over 13 million experiments on real-world data, for comprehensive evaluation of distributed optimization systems in the presence of adversaries.
  • A comprehensive evaluation of the adversarial impact on system efficiency and agent discomfort.
  • New insights highlighting adversarial influences on optimality in various application scenarios.
  • An open-source software artifact implementation of the proposed adversarial model for the I-EPOS collective learning algorithm [^1] Available at https://github.com/epournaras/EPOS.
The rest of this paper is organized as follows: Section II reviews related literature on adversarial impacts in distributed optimization. Section III introduces the proposed adversarial model, along with the problem formulation, detailing the network model and optimization challenge. Section IV outlines the experimental methodology and evaluation metrics. The key findings are presented in Section V, discussing system behavior under varying adversarial conditions. Finally, Section VI concludes the paper and outlines directions for future research.
II Related Work
Resilience in distributed multi-objective optimization plays a critical role across domains such as smart grids, transportation, logistics, and communication networks, where robust and adaptive systems are crucial for ensuring operational efficiency [21, 22]. Convex distributed optimization has received significant attention, with a focus on addressing challenges posed by adversary agents, network structures, and varied application domains [3, 13, 23]. Earlier work examines the robustness and vulnerability of consensus-based distributed optimization to address limitations and assumptions related to adversary behavior, network topology, objective functions, and application domains [13, 24, 25]. The presence of adversary agents significantly impacts the performance of distributed optimization models. These agents disrupt optimization by slowing convergence, manipulating data, or withholding participation, resulting in suboptimal performance [8, 15]. Table I provides a comparative analysis of related work in distributed optimization under adversarial conditions, focusing on adversary behavior, attack strategies, and their impact on network and system performance.
TABLE I: Comparison of literature on resilience in distributed optimization
Work Adversarial behavior Adversary target Knowledge of the system Network Performance measure Algorithms / Techniques for Optimization Recovery [10] Malicious & Byzantine  [^2] A node is considered malicious if it sends the same value to all its out-neighbors at each time-step, while a Byzantine node can send different values to different neighbors at each time-step. Consensus protocol Full knowledge of network structure and cost functions of other agents Connected undirected network Distance to optimality minimization Local filtering [26] Byzantine Consensus process & parameter agreement Full knowledge of network structure and cost functions of other agents Strongly connected directed network Convergence to a bounded region containing the global minimizer Distance-based filtering [27] Byzantine Information exchange among agents Full knowledge of network structure and cost functions of other agents Complete directed network Admissibility (structure of the convex coefficients) Local filtering [28] Adversarial manipulation of observations Single agent observation Black-box access to an agent’s policy and environment Limited directed communication Average reward & team win rate Gradient-based methods (Deep Q-learning) [29] Malicious Consensus-based MARL Partial knowledge (compromises rewards) Fully connected network Average team reward Consensus-based MARL [30] Data perturbations (cyber-attacks) DRL rewards Full knowledge of DRL model parameters DNN (power system topology) Expected performance decay & critical attack rate None [31] Injection & DoS attacks Consensus in cyber-physical systems Full knowledge of network structure and cost functions of neighbors Robust sparse network Resilient asymptotic consensus Mean subsequence reduced [32] Malicious data injections Convergence to the global optimal point Partial knowledge of network structure and neighbors’ states Undirected connected network Convergence to optimal point Probabilistic trust-based weight adjustment & projection-based update [33] Malicious & Byzantine Information exchange among agents Full knowledge of network structure and neighbors’ states Directed connected spanning network Convergence to optimal point Markov switching communication topology & Push-DIGing [34] Malicious & Byzantine Consensus process and state updates Full knowledge of network structure and neighbors’ states Undirected connected network Convergence to optimal point Resilient distributed optimization with trusted agents & connected dominating set [35, 36] Impersonation Byzantine Communication channels between agents and coordinator Local knowledge Directed network with central coordinator Distance to optimality minimization Primal-Dual [9] Malicious Consensus value Local knowledge Directed & strongly connected Distance to optimality minimization FROST [37] Malicious Consensus value Local knowledge Undirected network Distance to optimality minimization Randomized gossip [38, 39] Eavesdropping Information exchange among agents All transmitted messages Undirected network Convergence to a privacy-preserving approximate solution Differentially private gradient tracking This work Selfish malicious System-wide optimization Local knowledge & black-box access to network structure Bidirectional hierarchical network Optimization efficiency I-EPOS
II-A Adversary Agents in Distributed Optimization
Yang et al. [3] provide a comprehensive survey on distributed optimization. Notable advancements include extensions of consensus-based protocols by Sundaram et al. [10] and Kuwaranancharoen et al. [26], which address adversarial threats in convex optimization. Su et al. [27] enhance these methods with decentralized architectures and explore adversarial influence on global objectives. However, these approaches assume adversary agents have full knowledge of the network topology and the private functions of all agents. This coordination among adversaries compromises the privacy of the agents in the system.
II-B Adversarial Attacks in Multi-Agent Systems
Adversarial attacks significantly impact reinforcement learning (RL) systems across applications such as robotics, video games, and smart grids, undermining system stability and performance [40, 41]. Lin et al. [28] demonstrate how adversarial perturbations affect cooperative multi-agent RL (c-MARL), showing its vulnerability compared to single-agent RL. Figura et al. [29] highlight how a single adversary can influence consensus-based c-MARL systems, disrupting team objectives. Zheng et al. [30] introduce criticality-based perturbations in deep Q-networks, demonstrating substantial performance degradation due to adversarial attacks. These studies highlight that an adversary agent has the ability not only to disrupt system operations but also to manipulate the policy to align with its own adversarial objectives, thereby influencing other agents to adopt a policy of its preference. These insights underscore the critical need for robust defense against adversaries that manipulate policies and system states in RL systems.
II-C Cyber-Attacks and Resilient Control
Cyber-attacks, including data injection and denial-of-service (DoS) attacks, pose significant threats to distributed optimization by disrupting system operations and consensus mechanisms [31]. To address these challenges, Yemini et al. [32] introduce trust-based frameworks that mitigate malicious input, ensuring convergence to global optima. Similarly, Du et al. [33] and Zhao et al. [34] propose models relying on trusted agents to counteract adversarial influence. However, the effectiveness of these can be limited in scenarios with intermittent communication, such as ad hoc or robotic networks.
II-D Resource Allocation Challenges Under Adversaries
In resource allocation, adversarial disruptions are mitigated using robust optimization algorithms. Uribe et al. [35] and Turan et al. [36] propose primal-dual methods to tolerate adversaries by identifying and eliminating Byzantine inputs. Similarly, Ravi et al. [9] and Gentz et al. [37] focus on adversary detection, using statistical analysis and hypothesis testing to isolate malicious agents. These methods achieve resilience for up to 50% adversary density. However, they assume adversarial influence is evenly distributed and may not generalize to more concentrated or dynamic adversarial settings.
Another approach focuses on devising an adversary detection model to identify malicious agents in distributed optimization systems [9, 37]. Ravi et al. [9] address the challenge of detecting and mitigating malicious inputs in a distributed optimization system. Their method focuses on utilizing the data values of the agents to identify and isolate potential malicious agents. The researchers impose an upper limit on the number of tolerable malicious agents, specifically setting it at half of the total number of agents in the network topology. On the other hand, Gentz et al. propose a detection method for insider attackers in randomized gossip-based sensor networks. This approach involves utilizing theory and hypothesis testing to analyze the statistical states of the sensors, enabling the detection of malicious agents within the network [37].
II-E Privacy-Preserving Distributed Optimization
Privacy-preserving distributed optimization safeguards sensitive agent information against eavesdropping adversaries using techniques such as differential privacy [38, 39], homomorphic cryptography [42, 43], and gradient perturbation [44, 45], to ensure secure information exchange. However, these studies focus on privacy protection rather than adversarial behavior in optimization contexts.
II-F Combinatorial Optimization Algorithms
In distributed collective learning and combinatorial optimization, Hinrichs et al. [46, 47] propose COHDA, a combinatorial optimization heuristic designed for multi-agent systems. However, COHDA encounters scalability challenges due to increasing communication overhead as the number of agents grows. Pournaras et al. [48] address this with EPOS, a decentralized framework that enables agents to collaboratively optimize global resource allocation, particularly in participatory sharing economies. EPOS ensures privacy, autonomy, and scalability but encounters computational limitations when applied to wide tree structures with multiple child nodes [49]. I-EPOS extends EPOS by introducing decentralized iterative back-propagation and localized decision-making to enhance scalability [50, 49]. While these works significantly contribute to decentralized combinatorial optimization, none of them study the performance in the presence of adversary agents.
Research on adversarial attacks predominantly focuses on continuous distributed optimization, often analyzing specific attack types or assuming a limited number of adversaries, which may not accurately reflect real-world scenarios. These studies typically rely on complete graph assumptions and address single points of failure rather than system-wide resilience. While several solutions mitigate adversarial attacks in multi-agent optimization, there is a notable lack of comprehensive analyses of inherent system vulnerability, resilience, and how optimization collapses.
This study addresses the critical gap in understanding how adversary agents impact system functionality by introducing a generic adversarial model to analyze resilience and vulnerability in distributed multi-objective optimization. The primary objective is to assess in-depth the influence of adversary agents on system optimality. By filling this gap, this research aims to advance the design of self-healing strategies to enhance fault-tolerance and fault-correction in adversarial distributed optimization, which, to the best of our knowledge, has not been comprehensively explored in prior research.
III Adversarial Distributed Optimization
III-A Problem Formulation
Resilience in distributed optimization is essential for maintaining system performance under adversarial conditions. Adversary agents disrupt operations, degrade efficiency, and increase system vulnerability. This raises key questions: How do adversary agents influence the efficiency and stability of distributed optimization systems? What thresholds of adversarial behavior lead to transitions from resilience to vulnerability or collapse? How do parameters such as adversary density, adversarial severity, and network positioning influence these transitions?
To address these challenges, we propose a generic adversarial distributed optimization model tailored to discrete-choice scenarios. This model is used to investigate the trade-offs between system-wide and individual agent goals in adversarial settings. It incorporates critical parameters, including adversary density, adversarial severity, and structural positioning, providing a robust framework to evaluate system vulnerability and resilience. This study identifies thresholds for transitions from resilience to collapse, providing an in-depth understanding of system behavior under adversarial influence. These findings inform the development of self-healing strategies that enhance fault-tolerance and mitigate adversarial impacts across diverse distributed optimization applications.
III-B Network Model
Consider a network with $\mathsf{n}$ agents, denoted by $A$, each identified by a unique ID in the set $\{1,2,3,...,\mathsf{n}\}$. The network topology is represented as a connected graph $G=(A,E)$, where $A$ is the set of agents, and $E$ is the set of edges, such that $(j,i)\in E\neq(i,j)\in E$. Agents interact within a self-organized network through bidirectional communication to exchange information and update their states to align with the system goals.
Table II defines the notations used throughout the paper to formalize the network model and optimization framework.
TABLE II: Nomenclature utilized in this research
Notation Description
$A$ set of agents in the network
$\mathsf{n}$$=|A|$ number of agents
$A_{\mathsf{d}}\subseteq A$ set of adversary agents
$A_{\mathsf{l}}\subseteq A$ set of legitimate agents
$P_{\mathsf{a}}$ set of possible plans of agent $a$
[…]
$k$ number of plans
$d$ size of plan
$s_{\mathsf{a}}$ selected plan of agent $a$
$g$ global response
$D$ discomfort cost
$f_{D}$ discomfort cost function
$I$ inefficiency cost
$f_{I}$ inefficiency cost function
III-C Optimization Framework in Discrete-Choice Scenarios
[…]
Agents balance their individual preferences with system-wide goals, which often involve conflicting criteria. Each agent has an individual preference for its plans, calculated by the discomfort cost ($D$), such that $D$=${}_{a,i}=f_{D}(p_{\mathsf{a,i}})$, where $f_{D}$ is the agent’s preference associated with a particular plan. Each agent $a$ measures the costs $D$ for each possible plan $p_{\mathsf{a,i}}$$\in$ $P_{\mathsf{a}}$ and the plan with the minimum cost is selected. While agents aim to minimize their own discomfort, they may also consider system-wide metrics such as the inefficiency cost ($I$). The inefficiency cost ($I$) is the measure used to evaluate the collective system-wide performance based on the aggregated responses of all agents. It represents the system-wide performance inefficiency that agents aim to minimize through coordinated decision-making: $I=f_{I}\left(\sum_{a=1}^{\mathsf{n}}(s_{\mathsf{a}})\right)$.
Agents optimize their decisions by minimizing a weighted combination of these costs, as illustrated in Equation 1. The behavior of agent $a$ is modeled by the corresponding weights $\beta_{a},\text{ and }\alpha_{a}$, prioritizing efficiency and comfort, respectively. A higher weight indicates a higher preference for minimizing the corresponding objective. On the other hand, a weight of 0 means that the corresponding objective is not considered. For example, When $\beta_{a}=0,\text{ and }\alpha_{a}=1$, the agent prioritizes system-wide efficiency, acting as altruistic. Conversely, when $\beta_{a}=1,\text{ and }\alpha_{a}=0$, the agent prioritizes its own interest, behaving as selfish.
$s_{\mathsf{a}}$ $=\arg\min_{i=1}^{k}\left(\alpha_{a}\cdot I_{a,i}+\beta_{a}\cdot D_{a,i}\right)$
$=\arg\min_{i=1}^{k}\alpha_{a}\cdot f_{I}\left(s_{1},s_{2},\ldots,s_{\mathsf{n}}\right)$
$\quad+\beta_{a}\cdot f_{D}\left(D_{1,s},D_{2,s},\ldots,D_{\mathsf{n},s}\right)$ (1)
$\text{where}\quad\alpha_{a}+\beta_{a}$ $=1\quad\text{\&}\quad\alpha_{a},\beta_{a}\in[0,1]$
The decision-making process in distributed optimization is influenced not only by each agent’s individual objectives but also by the collective impact of their choices on system-wide goals. This interdependence introduces significant complexity, particularly when cost functions are non-linear, as agents’ choices depend on one another. Minimizing inefficiency and discomfort costs under such conditions is an NP-hard problem [49]. These challenges are further exacerbated by the lack of complete information about other agents’ selections and the need for coordination among agents’ choices. Addressing these issues requires a decentralized optimization framework that enables agents to iteratively refine their plan selections based on aggregated contributions and prior iterations (see Section IV-A).
III-D Adversarial Distributed Optimization Model
To study the resilience of distributed optimization in an adversarial environment, we partition the set of agents, $A$, into two subsets: adversary agents ($A_{\mathsf{d}}$) and legitimate agents ($A_{\mathsf{l}}$). Adversary agents, $A_{\mathsf{d}}\subseteq A$, prioritize their individual interests over collective system goals by adapting their behavior to maximize personal benefits. This behavior disregards system-wide efficiency and leads to suboptimal outcomes for the overall system. Legitimate agents, $A_{\mathsf{l}}\subseteq A$, align their actions with system-wide objectives to optimize overall performance. The total set of agents is defined as $A=A_{\mathsf{d}}\cup A_{\mathsf{l}}$.
Adversarial behavior amplifies the inefficiency cost $I$, reflecting the trade-off between minimizing individual discomfort ($D$) and optimizing overall system efficiency ($I$). While $D$ focuses on individual preferences, $I$ addresses the system-wide inefficiency caused by deviations from optimal resource allocation, underscoring the conflict between personal and collective goals. For instance, in a bike-sharing system, optimization ensures a balanced distribution of bikes across stations to meet user demand. Legitimate users select pick-up and drop-off stations with consideration for system-wide efficiency, aiming to maintain network equilibrium. Conversely, adversary users prioritize their convenience, choosing stations that minimize their discomfort without regard for the network state. This behavior creates imbalances, such as empty or overstocked stations, reducing overall efficiency and user satisfaction.
IV Experimental Methodology
This paper proposes a novel adversarial distributed optimization model to study the performance of multi-agent discrete-choice combinatorial optimization under distributed adversary attacks.
IV-A Decentralized Optimization Framework
The proposed model is implemented within the Iterative Economic Planning and Optimized Selections (I-EPOS) [^2] Available at https://github.com/epournaras/EPOS, a collective learning framework. I-EPOS employs a self-organized, multi-level hierarchical network structure that ensures efficient communication and coordinated decision-making while reducing communication overhead [51].
I-EPOS enables the agents to iteratively adjust their plan selections based on the aggregated plans of the agents beneath and previous iteration choices. The proposed model operates iteratively, enabling agents to coordinate their choices in collective decision-making. Each iteration consists of two distinct phases: a bottom-up phase and a top-down phase. During the bottom-up phase, agents select plans based on the aggregated choices of agents in the subtree beneath, as well as the selections made by all agents in the previous iteration. Conversely, the top-down phase addresses incomplete knowledge from higher branches in the hierarchy, enabling agents to revert to previous selections if no cost reduction is achieved. This process continues until a predefined iteration limit is reached or no further improvement in the optimization objective occurs [49, 50].
The hierarchical network structure is organized as an acyclic graph, where each parent agent aggregates responses from its children without redundancy, explicitly avoiding double counting of contributions from lower nodes in the hierarchy. This design ensures efficient coordination and alignment of individual decisions with system-wide objectives while maintaining overall optimization [49, 51, 52]. I-EPOS is well-suited for adversarial scenarios in Smart City applications due to its scalability, adaptability to diverse agent behaviors, and ability to handle adversarial conditions [50, 53].
IV-B Experimental Setup
Experiments are conducted using multiple HPC servers with varying configurations that support large-scale simulations and ensure computational efficiency. These include high-memory nodes (up to 768 GB) and multi-core processors (up to 40 cores per node). In addition to these servers, the University of Leeds ARC4 system  [^3] ARC4 is an HPC cluster at Leeds providing a Linux-based HPC service based on CentOS 7. More information: https://arcdocs.leeds.ac.uk/systems/arc4.html is utilized. The ARC4 system includes two nodes, each equipped with 40 cores, 768 GB of memory, and 800 GB of storage, providing robust computational capacity for large-scale experiments.
IV-C Application Scenarios
TABLE III: Description of the datasets and experimental setup in this research
Dataset Name No. Agents No. Plans Plan Size Agents Discomfort Cost Inefficiency Cost Total Experiments Energy 1000 10 144 Households Agent preference based on time shift Energy demand 30,018,000 Data-collective 72 3 64 participants Privacy loss Data collective inefficiency (quality of collected data) 65,433,600 (2 target signals) Voting 266 31 5 Voters Compromise Polarization 97,551,840 (120 target signals)
IV-C1 Energy-demand dataset
The energy application scenario uses a dataset derived from simulated zonal power transmission in the Pacific Northwest [^4] Available upon request at http://www.pnwsmartgrid.org/participants.asp. The dataset includes power consumption profiles for 1,000 users, with each user represented by an agent containing 10 possible plans. Each plan comprises a 144-length sequence representing electricity consumption at 5-minute intervals over a 12-hour period. These plans are generated using load-shifting strategies to balance grid load during peak and off-peak hours, reducing strain on the energy system. Plans are ranked by preference scores ranging from 0 to 1, with higher scores reflecting greater alignment with the user’s original consumption patterns. The inefficiency cost is measured as the deviation in aggregated energy consumption from the desired load-balancing levels, capturing the system ability to maintain stability and efficiency. Adversarial households disrupt the system by selecting plans that counteract load balancing, thereby increasing the risk of grid instability during peak periods.
IV-C2 Privacy dataset
The privacy dataset originates from a living-lab experiment at the Decision Science Laboratory [^5] https://www.descil.ethz.ch of ETH Zurich, involving 72 participants evaluating 64 data-sharing scenarios that involve 4 sensor types, data collectors, and contexts [53]. The data-sharing choices of each participant determine three data-sharing plans, representing their intrinsic motivation to share and two rewarded scenarios. The plans are assessed using privacy valuation schemes assigning normalized costs ranging from 0 to 1, where lower costs indicate less privacy compromise. The dataset facilitates testing under high and low privacy-preservation target signals, reflecting intrinsic preferences for data-sharing behaviors. The inefficiency cost is calculated based on the reduction in the quality of the aggregated data due to adversarial participants prioritizing minimal data sharing over collective goals. Adversarial participants disrupt coordination by focusing solely on minimizing their data sharing, under-mining the quality of service of data collectors.
IV-C3 Voting dataset
This new dataset is derived from voting data in a regional election with five candidates and 266 voters [^6] The UK Labour Party Leadership Vote Available at https://preflib.simonrey.fr/datasets [54]. Each voter has 31 alternative voting plans, representing ranked preferences among the five candidates. The optimization focuses on minimizing polarization in decision outcomes, which refers to the extent of divergence in voters’ preferences from a collectively agreed compromise. Inefficiency cost is measured as the cost of polarized decision outcomes. To ensure fair representation, 120 target signals are generated from all combinations of values 0, 0.25, 0.5, 0.75, and 1. Adversarial behavior occurs when voters prioritize minimizing their personal compromise cost, disregarding collective consensus, which amplifies polarization and disrupts the system ability to achieve shared, balanced outcomes.
Table III provides an overview of each dataset used in this research, including the number of agents, the number and size of plans per agent, agent representation, the corresponding discomfort and inefficiency cost functions, and the total number of experiments conducted.
IV-D Varying Dimensions and Performed Experiments
The experiments systematically evaluate the resilience and vulnerability of the system across the following dimensions:
  • Scale of adversaries ($|A_{\mathsf{d}}|$): Incrementally increase the number of adversary agents $A_{\mathsf{d}}$ from 1 to $\mathsf{n}$ across all datasets to analyze performance under varying adversary densities; i.e., $A_{\mathsf{d}}=\{a\mid a\in A\}$.
  • Adversarial severity ($\beta$): The adversarial preference to minimize discomfort cost ($\beta$) is varied across 30 levels, with $\beta$ incrementing from 0 to 1 such as $\beta=\frac{b}{30}$ for $b=\{1,2,3,...,30\}$. This variation captures the impact of varying levels of adversarial intent, where agents prioritize their personal objectives at the expense of system-wide goals, on overall system performance.
  • Adversary position: The influence of adversary positions within the hierarchical network is analyzed using two approaches: layer-wise and cumulative structural analysis. A binary tree structure is employed, with each hierarchical layer containing approximately $\log_{2}|A|$ agents, where $|A|$ is the total number of agents. The layer-wise analysis evaluates inefficiency costs across individual hierarchical layers under varying adversary scales (25%, 50%, 75%, 100%) and behavior levels ($\beta$ = 0.1, 0.4, 0.7, 1.0). The cumulative analysis examines the aggregated impact of adversary agents propagating through the network in top-down (root-to-leaf) and bottom-up (leaf-to-root) configurations.
[…]
IV-E Evaluation Metrics
Optimization objectives: Three metrics evaluate system performance: inefficiency cost $I$, discomfort cost $D$, and compromise cost of legitimate agents. The inefficiency cost measures the overall cost incurred by the system. The discomfort cost evaluates agents’ dissatisfaction with non-preferred choices. The compromise cost quantifies the extent to which legitimate agents adjust their individual preferences to align with collective goals under adversarial influence. This metric is calculated as the deviation between the selected plans of legitimate agents and their most preferred plans. These metrics evaluate the trade-offs between individual agent goals and system-wide objectives, particularly under varying adversarial conditions.
Pareto optimality: In multi-objective optimization, the Pareto front represents the set of non-dominated solutions where no objective can be improved without compromising another. The knee point on the Pareto front indicates the solution that achieves the best trade-off between competing objectives, making it a key decision-making criterion [55, 56].
This study identifies the knee point using the Minimum Manhattan Distance (MMD) method, which measures the distance between each solution on the Pareto front and an ideal reference point where both objectives are optimized. The solution with the smallest distance to the ideal point is selected as the knee point. This approach ensures a balanced evaluation of discomfort cost ($D$) and inefficiency cost ($I$) while maintaining alignment with established methodologies in the literature [56].
Resilience, vulnerability, and collapse framework: To define thresholds for resilience, vulnerability, and collapse in distributed optimization, we apply the multi-Otsu thresholding method. The Otsu method is an efficient technique for dividing data into distinct regions by minimizing intra-class variance [57, 58]. In this study, the method segments data into three regions based on inefficiency levels: resilience (low inefficiency), vulnerability (moderate inefficiency), and collapse (high inefficiency). This framework provides a robust method for segmenting inefficiency data into resilience, vulnerability, and collapse states, enabling a clear assessment of system performance transitions under adversarial conditions.
V Results Analysis and Discussion
This section presents the results of extensive experiments evaluating the impact of adversarial agents on multi-objective distributed optimization across the Energy, Voting, and Privacy datasets.
(a) Inefficiency costs
(b) Discomfort costs
(c) Compromised discomfort costs for non-adversary agents
Figure 1: Multi-objective optimization performance across different numbers of adversary agents and various levels of adversaries within the Energy, Voting, and Privacy datasets. R, V, and C indicate regions of Resilience, Vulnerability, and Collapse.
V-A Resilience Analysis
Figure 1 illustrates inefficiency costs, discomfort costs, and compromised discomfort costs across varying scale of adversaries and behavior levels. These metrics collectively capture system-wide inefficiency, individual agent dissatisfaction, and the trade-offs faced by legitimate agents under adversarial conditions.
Energy Dataset
[…]
Voting Dataset
The Voting dataset demonstrates significant stability, maintaining resilience across more than 90% of the operational space. Inefficiency remains low for $\beta<0.9$ and scale of adversaries below 50%. Vulnerability emerges between 50% and 70% adversary agents, with inefficiency increasing by 111%. Collapse occurs at $\beta\geq 0.96$ and proportions exceeding 70%. Discomfort costs decline gradually in resilient regions and are entirely eliminated under collapse conditions, highlighting a clear trade-off between individual comfort and system-wide efficiency. Compromised discomfort costs show resilience for adversary agents below 20%, remaining low even under high levels of adversarial behavior. However, collapse is observed for proportions exceeding 50% and adversarial levels $\beta>0.7$, where compromised discomfort costs rise smoothly. This trend suggests that legitimate agents in the Voting dataset show lower resilience compared to the Energy dataset.
Privacy Dataset
The Privacy dataset reveals lower resilience thresholds compared to the Energy and Voting datasets, especially for high privacy-preserving signal. For $\beta<0.2$ and adversary agents below 20%, inefficiency costs remain low. However, vulnerability emerges as proportions increase to 50%, and collapse occurs at $\beta\geq 0.3$ for high data-sharing levels and $\beta\geq 0.5$ for low data-sharing levels when adversary agents exceed 50%.
Discomfort costs follow similar trends, with resilience observed below 30% adversary agents for high privacy-preserving signal and 20% for low privacy-preserving signal. Collapse occurs when $\beta>0.3$ and proportions exceed 60%. Compromised discomfort costs rise steadily, peaking at values of 9 and 10 for high and low privacy-preserving signal, respectively, at 90% adversary agents. Unlike the Energy and Voting datasets, where $\beta$ strongly influences compromised discomfort costs, the Privacy dataset shows that scale of adversaries serves as the dominant driver of these costs beyond $\beta>0.3$. This highlights the Privacy dataset heightened sensitivity to adversarial density, requiring more targeted interventions to preserve legitimate agent objectives.
(a) Pareto front and knee points across adversarial severity.
(b) Pareto front and knee points across scales of adversaries.
Figure 2: The Pareto Optimality of The Energy Voting, and Privacy datasets
(a) Energy dataset.
(b) Voting dataset.
(c) Privacy dataset (High privacy-preserving).
(d) Privacy dataset (Low privacy-preserving).
Figure 3: Inefficiency costs across hierarchical structure layers under various adversarial configurations.
V-B Pareto Optimality Analysis
Pareto optimality analysis identifies Pareto fronts knee points that represent optimal system performance under varying adversarial conditions. Figure 2 highlights system resilience to adversarial severity and proportions across the Energy, Voting, and Privacy datasets. Figure 2(a) focuses on the impact of adversarial severity ($\beta$) on tolerated scale of adversaries, while Figure 2(b) explores how varying adversary densities affect tolerated behavior levels.
[…]
The Voting dataset demonstrates robust stability. Figure 2(a) shows consistent knee points tolerating 50–60% adversary agents across a broad range of behavior levels ($0.1\leq\beta\leq 0.7$). Even under high adversarial severity ($0.86\leq\beta\leq 0.9$), as illustrated in Figure 2(b), the system exhibits exceptional resilience, tolerating densities from 10% to 100%. This highlights the dataset high tolerance to adversarial influence.
The Privacy dataset exhibits slightly difference dynamics between high and low privacy-preserving signals. In Figure 2(a), tolerated scale of adversaries for high signal configurations are approximately 70% at $\beta<0.5$, but drop to 50% at higher intensities. The low signal Privacy scenario displays smoother declines, with tolerated densities averaging 68% Adversary agents across behavior levels. Figure 2(b) shows that high signal configurations begin at knee points of $\beta=0.03$ for 10% adversary agents, rising gradually to $\beta=0.2$ at 50–90%. Low privacy-preserving signal exhibits a smoother gradient, peaking at $\beta=0.36$ at 20% Adversary agents and declining to $\beta=0.13$ at full adversarial saturation.
V-C Structure Analysis
This section analyzes how hierarchical structures influence inefficiency costs in Energy, Voting, and Privacy datasets. This analysis is divided into two approaches: individual layer-wise and cumulative structure analysis.
V-C1 Layer-wise structural analysis
Figure 3 highlights inefficiency costs across hierarchical layers for varying scale of adversaries at 25%, 50%, 75%, and 100% and behavior levels ($\beta$) at 0.1, 0.4, 0.7, and 1.0. In all datasets, deeper layers and layers with higher agent proportions exhibit greater inefficiency escalation under intensified adversarial influence.
In the Energy dataset (Figure 3(a)), with 1,000 agents distributed across 10 layers, inefficiency costs remain stable up to Layer 5 at moderate adversarial severity ($\beta=0.1$, 0.4, and 0.7). Beyond Layer 6, costs fluctuate and peak at Layers 7 and 8. Notably, while Layer 9 contains the highest number of agents (256), inefficiency only peaks at $\beta=1.0$ and 100% Adversary agents, highlighting the compounded vulnerability of deeper layers under extreme adversarial influence. This emphasizes the increasing impact of both adversary density and adversarial severity on deeper layers.
The Voting dataset (Figure 3(b)), with 266 agents distributed across 9 layers, shows minimal inefficiency in the upper layers (1–5) across all adversarial configurations. However, inefficiency rises sharply beyond Layer 5, with Layer 7 consistently incurring the highest costs, especially at higher adversarial severity ($\beta\geq 0.7$) and scale of adversaries of 75% and 100%. At $\beta=1.0$, inefficiency increases steadily from the root (Layer 1) to the leaves (Layer 9), highlighting the compounding impact of adversarial behavior as it propagates deeper into the hierarchy.
The Privacy dataset (Figures 3(c) and 3(d)), with 72 agents distributed across 7 layers, exhibits almost similar trends for high and low privacy-preserving signal. At low adversarial behavior ($\beta=0.1$), inefficiency costs remain stable across all layers, demonstrating resilience to mild adversarial influence. At higher levels of adversarial behavior ($\beta\geq 0.7$), inefficiency peaks at Layer 6, which has the highest agent proportion (32 agents). Interestingly, Layer 7, with fewer agents, incurs lower costs even under severe adversarial conditions. Minor deviations, such as Layer 5 occasionally surpassing Layer 6 in the low signal scenario, highlight the nuanced relationship between layer depth, agent density, and adversarial behavior.
These results suggest that inefficiency costs in hierarchical systems are highly sensitive to layer depth, scale of adversaries, and the level of adversarial behavior. Deeper layers and those with higher agent proportion emerge as critical vulnerabilities under intensified adversarial configurations. Reinforcing deeper layers and redistributing agent densities could mitigate inefficiency spikes, especially in applications with high adversarial risks.
(a) Top-down cumulative
(b) Bottom-up cumulative
Figure 4: Inefficiency cost across hierarchical structure layers under various adversarial configurations in Energy, voting and Privacy datasets.
V-C2 Cumulative structural analysis
Another approach assesses the cumulative impact of adversary agents on inefficiency costs as they propagate through the hierarchical network structure in two configurations: top-down (root-to-leaf) and bottom-up (leaf-to-root). Figure 4 presents inefficiency trends for scale of adversaries and adversarial severity (0.1, 0.4, 0.7, 1.0) across Energy, Voting, and Privacy datasets.
The cumulative structural analysis of the Energy dataset reveals distinct differences in inefficiency trends between top-down and bottom-up propagation. At $\beta=1$, top-down propagation results in significantly higher initial inefficiency costs, reaching 272 with a single adversarial agent, compared to just 0.8 in bottom-up propagation. Inefficiency costs in the top-down configuration also rise more rapidly, indicating accelerated system deterioration. In contrast, bottom-up propagation exhibits a slower emergence of inefficiency, with costs remaining low up to 20% adversarial agent participation, followed by exponential growth. These trends emphasize the critical impact of adversarial agent placement, as costs escalate faster and peak higher when adversaries are positioned in top layers. At lower adversarial severity (0.1, 0.4, 0.7), inefficiency costs remain low and stable, fluctuating between 0.1 and 0.5, even as number of adversaries increases. This highlights the system resilience to moderate adversarial behavior.
The Voting dataset shows distinct trends in inefficiency costs. At $\beta=1$, inefficiency costs start at approximately 2.8, similar to lower $\beta$ levels, but rise sharply as the proportion of adversary agents increases, peaking at 6.0 when adversary agents saturate the network (100%). In the top-down propagation, inefficiency costs accelerate rapidly within the top four layers, driven by early adversarial influence. This is followed by a clear decrease between Layers 5 and 6, corresponding to scale of adversaries ranging from 6% to 17%, indicating that inefficiencies stabilize as the system progresses deeper into the hierarchy. At lower adversarial severity ($\beta=0.1,0.4,0.7$), inefficiency costs remain low and stable, fluctuating slightly between 2.8 and 2.9, regardless of scale of adversaries. Notably, at $\beta=0.7$, inefficiency costs increase faster in the top-down propagation compared to the bottom-up propagation, underscoring the impact of adversary agents propagation direction on system performance under moderate adversarial behavior.
The Privacy dataset reveals resilience at low adversarial severity ($\beta=0.1$), with minimal inefficiency increases (0.01–0.05) across all layers in both top-down and bottom-up propagations, even under full adversary participation. At higher adversarial levels ($\beta=0.7,0.4$ and $\beta=1$), inefficiency costs rise significantly across layers. For high privacy-preserving signal, inefficiency emerges early with just one adversary agent in top-down propagation at Layer 1 with $\beta=0.4,0.7$, and $1$, while in bottom-up propagation, inefficiency begins with low value $<62.8$. In both propagations Layer 2 shows heightened sensitivity, where even 2-3 adversary agents sharply increase inefficiency. Layer 6 experiences notable surges when the scale of adversaries exceeds 50% for both propagation directions. In low privacy-preserving signals, inefficiency costs rise more prominently in top-down propagation, particularly at Layer 5, with sharp increases at $\beta=0.4,0.7$, and $1$ when scale of adversaries exceeds 25%. In bottom-up propagation, inefficiency escalates at $\beta=0.7$ and $1$ when scale of adversaries exceeds 70%. These patterns emphasize structural vulnerabilities, with Layers 2, 5, and 6 emerging as critical points for inefficiency escalation under intensified adversarial influence.
V-D Discussion
The results of this study reveal critical insights into adversarial influence on multi-objective distributed optimization systems, uncovering opportunities for designing self-healing strategies that enhance system resilience. By examining the dynamics of resilience, vulnerability, and collapse across diverse datasets, the analysis offers valuable lessons for addressing challenges in adversarial environments.
The Energy and Voting datasets exhibit strong resilience, withstanding a broad range of adversarial configurations and transitioning smoothly from vulnerability to collapse under higher adversarial severity. In contrast, the Privacy dataset demonstrates heightened sensitivity to adversarial scales and severity, particularly under high privacy-preservation constraints. This results underscore the importance of tailoring optimization strategies, particularly for privacy-preserving systems.
Pareto analysis identifies critical resilience thresholds where systems can balance inefficiency and discomfort costs. The Energy dataset sustains resilience up to 90% adversarial scale, collapsing only under extreme severity, while the Voting dataset maintains consistent performance across configurations, showcasing robustness. The Privacy dataset requires earlier interventions for high privacy-preserving configurations but allows for more gradual adjustments under low privacy-preserving scenarios. These findings reinforce the importance of identifying Pareto knee points to enable dynamic trade-off prioritization and effective system adaptation under real-time adversarial conditions.
Structural analysis reveals vulnerabilities in deeper hierarchical layers with higher number of agents, which are more prone to inefficiency propagation under intensified adversarial influence. Targeted strategies, such as reinforcing critical layers, redistributing agent densities, and employing localized containment mechanisms, can mitigate inefficiencies. Top-down configurations show rapid cascading of adversarial impact from top layers, demanding robust protections at higher levels, whereas bottom-up configurations exhibit slower propagation, suggesting inherent resilience.
Integrating adaptive fault-monitoring mechanisms is essential for maintaining system stability. Dynamically detecting transitions between resilience, vulnerability, and collapse allows for strategic resource allocation. For instance, energy systems can employ load-balancing strategies to address inefficiency, while hierarchical vulnerabilities in distributed optimization algorithms should be addressed through structural reinforcements and proactive monitoring.
This research provides a comprehensive framework for understanding adversarial impacts on distributed optimization systems and offers actionable insights for designing robust, self-healing mechanisms. By combining structural interventions, adaptive monitoring, and informed resource allocation, distributed systems can maintain efficiency and resilience even under significant adversarial configurations.
VI Conclusion and Future work
This study provides a comprehensive analysis of resilience, vulnerability, and collapse dynamics in multi-agent distributed optimization systems under adversarial conditions. By systematically examining the impact of the scale of adversaries, adversarial severity, and hierarchical structure, it identifies critical thresholds where system performance transitions from resilience to collapse. These insights offer practical strategies to enhance system design and resilience against adversarial influences.
The results highlight critical thresholds where systems transition from resilience to collapse, emphasizing the interplay of structural vulnerabilities and adversarial factors. These findings provide a foundation for designing tailored resilience strategies across diverse operational contexts.
A significant contribution of this work is the introduction of a benchmark dataset and software artifact for I-EPOS collective learning. These resources enable researchers to further investigate adversarial dynamics and develop advanced self-healing strategies. Integrating real-time detection and mitigation mechanisms will be critical for dynamically restoring system performance and resilience in the presence of adversarial behaviors.
Future work will focus on embedding real-time adaptive mechanisms within practical systems to monitor and counteract adversarial influences dynamically. Exploring complex network topologies, diverse adversarial scenarios, and broader application domains will deepen understanding and support the development of robust, fault-tolerant distributed optimization systems capable of maintaining efficiency and stability under challenging conditions.
Acknowledgments
This project is funded by a UKRI Future Leaders Fellowship (MR-/W009560-/1): ‘Digitally Assisted Collective Governance of Smart City Commons–ARTIO’. The authors would like to thank Thomas Wellings and Chuhao Qin for their feedback on the paper, Abhinav Sharma for technical support in experimentation and Srijoni Majumdar for support with the voting dataset.

### v2

@@
 Abstract
-Optimization is instrumental for improving operations of large-scale socio-technical infrastructures of Smart Cities, for instance, energy and traffic systems.
-In particular, understanding the performance of multi-agent discrete-choice combinatorial optimization under distributed adversary attacks is a compelling and underexplored problem, since multi-agent systems exhibit a large number of remote control variables that can influence in an unprecedented way the cost-effectiveness of distributed optimization heuristics.
-This paper unravels for the first time the trajectories of distributed optimization from resilience to vulnerability, and finally to collapse under varying adversary influence.
-Using real-world data to emulate over 28 billion multi-agent optimization scenarios, we exhaustively assess how the number of agents with different adversarial severity and network positioning influences optimization performance, including the influence on Pareto optimal points.
-With this novel large-scale dataset, made openly available as a benchmark, we disentangle how optimization remains resilient to adversaries and which adversary conditions are required to make optimization vulnerable or collapsed.
-These new findings can provide new insights for designing self-healing strategies for fault-tolerance and fault-correction in adversarial distributed optimization that have been missing so far.
+Optimization is critical for improving the operations of large-scale socio-technical infrastructures such as those found in energy, mobility, and information systems.
+In particular, understanding the performance of multi-agent discrete-choice combinatorial optimization under distributed adversarial attacks is a compelling and underexplored problem.
+Multi-agent systems involve a large number of remote control variables that can influence the cost-effectiveness of distributed optimization heuristics.
+This paper unravels, for the first time, the trajectories of distributed optimization from resilience to vulnerability, and finally to collapse under varying adversarial influence.
+Using real-world and synthetic data to generate over 112 million multi-agent optimization scenarios, we systematically assess how the number of agents with varying levels of adversarial severity and network positioning influences optimization performance, with particular attention to the impact on Pareto optimality.
+With this large-scale dataset, made openly available as a benchmark, we disentangle how optimization systems remain resilient to adversaries and which adversary conditions make optimization vulnerable or cause collapse.
+These findings can support the design of self-healing strategies for fault tolerance and fault correction, addressing a critical gap in adversarial distributed optimization.
 Index Terms:
@@
 I Introduction
-The rapid development of Internet of Things (IoT) applications has brought about transformative changes in numerous domains, ranging from smart cities to industrial automation and healthcare [1].
-These applications involve vast networks of interconnected devices that generate substantial amounts of data, and require efficient and decentralized decision-making for processing [2].
+The rapid development of Internet of Things (IoT) applications has brought transformative changes in numerous domains, ranging from smart cities to industrial automation and healthcare [1].
+These applications involve vast networks of interconnected devices that generate substantial amounts of data, and require efficient and distributed decision-making [2].
 Distributed optimization is essential in such scenarios, as it enables multiple agents to collaborate effectively without centralized coordination.
@@
 While distributed optimization is essential for achieving system-wide objectives, most algorithms rely on assumptions of rational, cooperative, and non-adversarial agents contributing toward the global objective without prioritizing their individual goals over collective outcomes.
-However, real-world scenarios often deviate from these assumptions, as adversarial agents can disrupt optimization systems by introducing inaccuracies, manipulating decisions, or compromising functionality [8, 9, 10, 11].
-For example, in smart grid systems, where autonomous agents (e.g., households) collaborate to manage energy distribution and prevent power outages, certain households may act adversarially by manipulating consumption data and prioritizing self-interest, i.e.
-their thermal comfort.
-Such disruptions distort energy allocation processes, leading to inefficient resource use and blackouts [9, 10, 12, 13, 14].
+However, real-world scenarios often deviate from these assumptions; adversary agents can disrupt optimization by introducing inaccuracies, manipulating decisions, or compromising functionality [8, 9, 10, 11].
+For example, in smart grid systems where autonomous agents (e.g., households) collaborate to manage energy distribution and prevent power outages, certain households may act adversarially by manipulating consumption data to prioritize their self-interest, i.e., their thermal comfort.
+Such disruptions distort energy allocation, leading to inefficient resource use and blackouts [9, 10, 12, 13, 14].
 Several studies have examined multi-agent optimization in continuous-choice frameworks [8, 9, 10, 11, 13, 15].
 While stochastic optimization techniques address noise, they lack mechanisms to counteract strategic manipulations by adversarial agents [16, 17].
-These manipulations amplify system vulnerability by prioritizing individual goals over collective objectives, which can lead to inefficiencies, instability, and eventual collapse [18, 19].
-Despite this, adversarial disruptions in discrete-choice settings have received little attention.
-This paper addresses this gap by investigating the resilience, vulnerability, and collapse of multi-agent systems in discrete-choice optimization under adversarial conditions, offering a novel framework for understanding system behavior.
-Resilience refers to the system ability to maintain performance despite adversarial influence.
-Vulnerability captures the emergence of inefficiencies due to adversarial behaviors, and collapse occurs when the system fails to achieve its objectives.
-This study evaluates the impact of adversarial agents on optimization performance, analyzing critical parameters such as the number of adversarial agents, their adversarial severity, and their network positions.
-By systematically examining these dynamics, the paper provides critical insights into the conditions under which systems transition from resilience to collapse, unravelling new insights for the development of strategies [20] for self-healing, fault-tolerance, and fault-correction in adversarial environments.
+Such manipulations amplify system vulnerability by prioritizing individual goals over collective objectives, which can lead to inefficiencies, instability, and eventual system collapse [18, 19].
+Despite these risks, adversarial disruptions in discrete-choice settings have received little attention.
+This paper studies multi-agent systems in discrete-choice combinatorial optimization under adversarial conditions, with the aim to unravel the trajectories of resilience, vulnerability, and collapse, offering a novel framework to understand system optimization behavior.
+In this context, resilience refers to the system ability to maintain performance despite adversarial influence; vulnerability captures the emergence of inefficiencies due to intolerable adversarial behaviors; and collapse occurs when the system significantly under-performs as a result of failure to cope with adversarial behaviors.
+This study evaluates the impact of adversarial agents on optimization performance, analyzing critical parameters such as the number of adversarial agents, their behavioral severity, and their network positions.
+By systematically examining these dynamics, the paper provides critical insights into the conditions under which systems transition from resilience to collapse, offering new insights for developing self-healing, fault-tolerant, and fault-correcting strategies in adversarial environments [20].
 The main contributions of this paper are outlined as follows:
   • An adversarial model for discrete-choice multi-objective optimization problems.
-  • A novel evaluation framework for characterizing resilience, vulnerability, and collapse in distributed optimization systems under adversarial conditions.
-  • The applicability of the proposed adversarial model in different application domains.
-  • Open large-scale benchmark datasets comprising over 13 million experiments on real-world data, for comprehensive evaluation of distributed optimization systems in the presence of adversaries.
+  • A novel evaluation framework for characterizing resilience, vulnerability, and collapse in distributed optimization systems under adversarial influence.
   • A comprehensive evaluation of the adversarial impact on system efficiency and agent discomfort.
-  • New insights highlighting adversarial influences on optimality in various application scenarios.
-  • An open-source software artifact implementation of the proposed adversarial model for the I-EPOS collective learning algorithm [^1] Available at https://github.com/epournaras/EPOS.
-The rest of this paper is organized as follows: Section II reviews related literature on adversarial impacts in distributed optimization.
-Section III introduces the proposed adversarial model, along with the problem formulation, detailing the network model and optimization challenge.
+  • The first open large-scale benchmark datasets for discrete-choice adversarial optimization, generated from over 112 million experiments on real-world and synthetic inputs using the proposed adversarial model.
+  • New insights into how adversarial scale, severity, and structural positioning affect system optimality, including resilience and vulnerability thresholds, Pareto trade-offs, and structural vulnerabilities across diverse scenarios.
+  • An open-source software artifact implementing the proposed adversarial model within the I-EPOS system [^1] Available at https://github.com/epournaras/EPOS, extending its functionality to support heterogeneous agent behaviors.
+The rest of this paper is organized as follows: Section II reviews related work on adversarial distributed optimization.
+Section III introduces the proposed adversarial model and problem formulation, including the network model and optimization challenge.
 Section IV outlines the experimental methodology and evaluation metrics.
-The key findings are presented in Section V, discussing system behavior under varying adversarial conditions.
-Finally, Section VI concludes the paper and outlines directions for future research.
+Section V presents the key findings, analyzing system behavior under varying adversarial conditions.
+Finally, Section VI concludes the paper, discusses its limitations, and outlines directions for future research.
 II Related Work
@@
 Convex distributed optimization has received significant attention, with a focus on addressing challenges posed by adversary agents, network structures, and varied application domains [3, 13, 23].
-Earlier work examines the robustness and vulnerability of consensus-based distributed optimization to address limitations and assumptions related to adversary behavior, network topology, objective functions, and application domains [13, 24, 25].
+Earlier work examined the robustness and vulnerability of consensus-based distributed optimization, focusing on addressing limitations related to adversary behavior, network topology, objective functions, and application domains [13, 24, 25].
 The presence of adversary agents significantly impacts the performance of distributed optimization models.
 These agents disrupt optimization by slowing convergence, manipulating data, or withholding participation, resulting in suboptimal performance [8, 15].
-Table I provides a comparative analysis of related work in distributed optimization under adversarial conditions, focusing on adversary behavior, attack strategies, and their impact on network and system performance.
+Table I provides a comparative analysis of related work on distributed optimization under adversarial conditions.
+It highlights key aspects such as the type of adversarial behavior [^2] A malicious node sends the same value to all its neighbors at each time step, whereas a Byzantine node may send different values to different neighbors., attack targets, system knowledge [^3] Knowledge levels: full—complete knowledge of the network and agent objectives; partial—access to limited neighbor information; local—only own state or data is known., network structures, and the impact on overall performance.
 TABLE I: Comparison of literature on resilience in distributed optimization
-Work Adversarial behavior Adversary target Knowledge of the system Network Performance measure Algorithms / Techniques for Optimization Recovery [10] Malicious & Byzantine  [^2] A node is considered malicious if it sends the same value to all its out-neighbors at each time-step, while a Byzantine node can send different values to different neighbors at each time-step.
-Consensus protocol Full knowledge of network structure and cost functions of other agents Connected undirected network Distance to optimality minimization Local filtering [26] Byzantine Consensus process & parameter agreement Full knowledge of network structure and cost functions of other agents Strongly connected directed network Convergence to a bounded region containing the global minimizer Distance-based filtering [27] Byzantine Information exchange among agents Full knowledge of network structure and cost functions of other agents Complete directed network Admissibility (structure of the convex coefficients) Local filtering [28] Adversarial manipulation of observations Single agent observation Black-box access to an agent’s policy and environment Limited directed communication Average reward & team win rate Gradient-based methods (Deep Q-learning) [29] Malicious Consensus-based MARL Partial knowledge (compromises rewards) Fully connected network Average team reward Consensus-based MARL [30] Data perturbations (cyber-attacks) DRL rewards Full knowledge of DRL model parameters DNN (power system topology) Expected performance decay & critical attack rate None [31] Injection & DoS attacks Consensus in cyber-physical systems Full knowledge of network structure and cost functions of neighbors Robust sparse network Resilient asymptotic consensus Mean subsequence reduced [32] Malicious data injections Convergence to the global optimal point Partial knowledge of network structure and neighbors’ states Undirected connected network Convergence to optimal point Probabilistic trust-based weight adjustment & projection-based update [33] Malicious & Byzantine Information exchange among agents Full knowledge of network structure and neighbors’ states Directed connected spanning network Convergence to optimal point Markov switching communication topology & Push-DIGing [34] Malicious & Byzantine Consensus process and state updates Full knowledge of network structure and neighbors’ states Undirected connected network Convergence to optimal point Resilient distributed optimization with trusted agents & connected dominating set [35, 36] Impersonation Byzantine Communication channels between agents and coordinator Local knowledge Directed network with central coordinator Distance to optimality minimization Primal-Dual [9] Malicious Consensus value Local knowledge Directed & strongly connected Distance to optimality minimization FROST [37] Malicious Consensus value Local knowledge Undirected network Distance to optimality minimization Randomized gossip [38, 39] Eavesdropping Information exchange among agents All transmitted messages Undirected network Convergence to a privacy-preserving approximate solution Differentially private gradient tracking This work Selfish malicious System-wide optimization Local knowledge & black-box access to network structure Bidirectional hierarchical network Optimization efficiency I-EPOS
+[10] [26] [27] [28] [29] [30] [31] [32] [33] [34] [35, 36] [9] [37] [38, 39] This work Adversarial behavior Malicious Byzantine Cyber‐attacks Eavesdropping Adversary target Consensus Information Exchange System Objective Observation Knowledge of the system Full Partial Local Directed Network Performance measure Convergence Distance to Optimality Reward/Utility Efficiency Algorithms / Techniques for Optimization Local filtering Distance– based filtering Local filtering Gradient –based (Deep Q‐learning) Consensus‐based MARL None Mean subsequence reduced Probabilistic trust– based & projection–based Markov switching topology & Push‐ DIGing Resilience with trusted agents & dominating set Primal– Dual FROST Randomized gossip Differentially private gradient tracking I‐EPOS
 II-A Adversary Agents in Distributed Optimization
@@
 [30] introduce criticality-based perturbations in deep Q-networks, demonstrating substantial performance degradation due to adversarial attacks.
-These studies highlight that an adversary agent has the ability not only to disrupt system operations but also to manipulate the policy to align with its own adversarial objectives, thereby influencing other agents to adopt a policy of its preference.
-These insights underscore the critical need for robust defense against adversaries that manipulate policies and system states in RL systems.
+These studies show that an adversary can disrupt system operations and manipulate policies, influencing other agents to adopt behaviors aligned with its objectives.
 II-C Cyber-Attacks and Resilient Control
@@
 II-D Resource Allocation Challenges Under Adversaries
-In resource allocation, adversarial disruptions are mitigated using robust optimization algorithms.
+In distributed resource allocation, adversarial disruptions are typically mitigated using robust optimization and detection mechanisms.
 Uribe et al.
 [35] and Turan et al.
-[36] propose primal-dual methods to tolerate adversaries by identifying and eliminating Byzantine inputs.
+[36] propose primal-dual methods that tolerate Byzantine adversaries by identifying and eliminating malicious inputs, achieving resilience for up to 50% adversary density.
 Similarly, Ravi et al.
-[9] and Gentz et al.
-[37] focus on adversary detection, using statistical analysis and hypothesis testing to isolate malicious agents.
-These methods achieve resilience for up to 50% adversary density.
-However, they assume adversarial influence is evenly distributed and may not generalize to more concentrated or dynamic adversarial settings.
-Another approach focuses on devising an adversary detection model to identify malicious agents in distributed optimization systems [9, 37].
-Ravi et al.
-[9] address the challenge of detecting and mitigating malicious inputs in a distributed optimization system.
-Their method focuses on utilizing the data values of the agents to identify and isolate potential malicious agents.
-The researchers impose an upper limit on the number of tolerable malicious agents, specifically setting it at half of the total number of agents in the network topology.
-On the other hand, Gentz et al.
-propose a detection method for insider attackers in randomized gossip-based sensor networks.
-This approach involves utilizing theory and hypothesis testing to analyze the statistical states of the sensors, enabling the detection of malicious agents within the network [37].
-II-E Privacy-Preserving Distributed Optimization
-Privacy-preserving distributed optimization safeguards sensitive agent information against eavesdropping adversaries using techniques such as differential privacy [38, 39], homomorphic cryptography [42, 43], and gradient perturbation [44, 45], to ensure secure information exchange.
+[9] develop a detection method that uses agents’ data values to identify and isolate potential malicious behavior, imposing an upper limit of 50% adversaries within the network topology.
+Gentz et al.
+[37] propose a detection method based on hypothesis-testing for insider attackers in randomized gossip-based sensor networks, leveraging statistical analysis of sensor states to identify malicious agents.
+While these methods enhance resilience against dispersed adversaries, they assume adversarial influence is evenly distributed and may not generalize to scenarios with concentrated or dynamic adversary placement.
+II-E Multi-Objective Distributed Optimization
+Recent studies have increasingly adopted Pareto-based multi-objective optimization to evaluate system trade-offs in complex infrastructures.
+Fettah et al.
+[42] introduce a Pareto strategy for optimizing distributed generation in power networks.
+Zhang et al.
+[43] formulate a multi-objective operational framework that integrates Pareto analysis to enhance resilience thresholds in distribution networks.
+Similarly, Boindala and Ostfeld [44] propose an optimization approach to balance reliability, cost, and failure risk using Pareto fronts.
+While these studies underscore the value of Pareto analysis for resilient optimization, they focus on centralized infrastructures and do not address adversarial influence or the complexities of distributed, multi-agent decision-making explored in this work.
+II-F Privacy-Preserving Distributed Optimization
+Privacy-preserving distributed optimization safeguards sensitive agent information against eavesdropping adversaries using techniques such as differential privacy [45, 38, 39], homomorphic cryptography [46, 47], and gradient perturbation [48, 49], to ensure secure information exchange.
 However, these studies focus on privacy protection rather than adversarial behavior in optimization contexts.
-II-F Combinatorial Optimization Algorithms
-In distributed collective learning and combinatorial optimization, Hinrichs et al.
+II-G Combinatorial Optimization Algorithms
+In distributed combinatorial optimization, Hinrichs et al.
 [50, 51] propose COHDA, a combinatorial optimization heuristic designed for multi-agent systems.
 However, COHDA encounters scalability challenges due to increasing communication overhead as the number of agents grows.
-Pournaras et al.
-[48] address this with EPOS, a decentralized framework that enables agents to collaboratively optimize global resource allocation, particularly in participatory sharing economies.
-EPOS ensures privacy, autonomy, and scalability but encounters computational limitations when applied to wide tree structures with multiple child nodes [49].
-I-EPOS extends EPOS by introducing decentralized iterative back-propagation and localized decision-making to enhance scalability [50, 49].
-While these works significantly contribute to decentralized combinatorial optimization, none of them study the performance in the presence of adversary agents.
-Research on adversarial attacks predominantly focuses on continuous distributed optimization, often analyzing specific attack types or assuming a limited number of adversaries, which may not accurately reflect real-world scenarios.
-These studies typically rely on complete graph assumptions and address single points of failure rather than system-wide resilience.
-While several solutions mitigate adversarial attacks in multi-agent optimization, there is a notable lack of comprehensive analyses of inherent system vulnerability, resilience, and how optimization collapses.
-This study addresses the critical gap in understanding how adversary agents impact system functionality by introducing a generic adversarial model to analyze resilience and vulnerability in distributed multi-objective optimization.
-The primary objective is to assess in-depth the influence of adversary agents on system optimality.
-By filling this gap, this research aims to advance the design of self-healing strategies to enhance fault-tolerance and fault-correction in adversarial distributed optimization, which, to the best of our knowledge, has not been comprehensively explored in prior research.
+The collective learning approach of Pournaras et al.
+[52] address this with EPOS (Economic Planning and Optimized Selections), a distributed optimization method that enables agents to collaboratively optimize global resource allocation, particularly in participatory sharing economies.
+Although EPOS ensures privacy, autonomy, and scalability, it faces computational limitations when applied to wide tree structures with multiple child nodes [53].
+I-EPOS is the iterative extension of EPOS; it incorporates decentralized iterative back-propagation and localized decision-making to enhance scalability and support plan coordination across deeper and broader network hierarchies [54, 53].
+While COHDA, EPOS, and I-EPOS represent foundational combinatorial optimization approaches, their system performance under adversarial conditions has not been studied before, despite some limited work on measuring the impact of arbitrary structural faults, which does not focus on agents’ behavior [55].
+In this work, we introduce a generic adversarial model applicable to such settings.
+The model enables a structured evaluation of resilience, vulnerability, and collapse dynamics, and, to the best of our knowledge, is the first to systematically explore adversarial behaviors in discrete-choice combinatorial optimization.
+A large body of research to date focuses on continuous distributed optimization, often assuming limited adversary presence or relying on complete graph topologies [8, 9, 10, 13].
+Such assumptions do not fully capture the complexity of real-world systems, where distributed structures, heterogeneous agent behaviors, and dynamic adversarial threats are prevalent [30, 10].
+Although existing solutions offer valuable mitigation strategies [35, 33], the lack of comprehensive analyses on inherent system vulnerability, resilience thresholds, and pathways to optimization collapse remains a gap.
+To fill this gap, we propose a generic adversarial model to systematically analyze how adversarial agents influence system performance and stability in distributed multi-objective optimization.Our objective is to evaluate how adversarial behavior influences resilience, vulnerability, and collapse, while informing the development of self-healing strategies for robust optimization.
+Moreover, his work releases the first large-scale, open benchmark dataset designed for evaluating adversarial impacts under discrete-choice optimization settings.
+To clarify the novelty of our contribution, Table II summarizes a comparative analysis with existing work, highlighting key aspects such as discrete decision-making, multi-objective formulation, resilience thresholds, structural vulnerability, and Pareto-based evaluation.
+TABLE II: Comparison of the novelty aspects of this work with related distributed optimization approaches
+Criteria This work [10] [26] [27] [28] [29] [30] [31] [32, 33, 34] [35, 36] [9] [37] [38, 39] [42, 43, 44] Multi-objective optimization X X X X X X Discrete decision-making X X X X X X X X X X Resilience & vulnerability thresholds X X X X Structure analysis X X X X X X X X X Pareto analysis X X X X X X X X X X X X
 III Adversarial Distributed Optimization
@@
 Resilience in distributed optimization is essential for maintaining system performance under adversarial conditions.
-Adversary agents disrupt operations, degrade efficiency, and increase system vulnerability.
+Adversary agents disrupt operations, degrade efficiency, and increase vulnerability.
 This raises key questions: How do adversary agents influence the efficiency and stability of distributed optimization systems?
@@
 To address these challenges, we propose a generic adversarial distributed optimization model tailored to discrete-choice scenarios.
-This model is used to investigate the trade-offs between system-wide and individual agent goals in adversarial settings.
-It incorporates critical parameters, including adversary density, adversarial severity, and structural positioning, providing a robust framework to evaluate system vulnerability and resilience.
-This study identifies thresholds for transitions from resilience to collapse, providing an in-depth understanding of system behavior under adversarial influence.
+Our model investigates the trade-offs between system-wide and individual agent goals in adversarial settings.
+It incorporates key parameters, including the scale of adversaries, behavioral severity, and structural positioning, providing a robust framework to evaluate system vulnerability and resilience.
+Through this study, we identify critical thresholds for transitions from resilience to collapse, providing an in-depth understanding of system behavior under adversarial influence.
 These findings inform the development of self-healing strategies that enhance fault-tolerance and mitigate adversarial impacts across diverse distributed optimization applications.
@@
 Consider a network with $\mathsf{n}$ agents, denoted by $A$, each identified by a unique ID in the set $\{1,2,3,...,\mathsf{n}\}$.
-The network topology is represented as a connected graph $G=(A,E)$, where $A$ is the set of agents, and $E$ is the set of edges, such that $(j,i)\in E\neq(i,j)\in E$.
-Agents interact within a self-organized network through bidirectional communication to exchange information and update their states to align with the system goals.
-Table II defines the notations used throughout the paper to formalize the network model and optimization framework.
-TABLE II: Nomenclature utilized in this research
+The network topology is represented as a connected graph $G=(A,E)$, where $A$ is the set of agents, and $E$ is the set of edges, with $(j,i)\in E$ not necessarily implying $(i,j)\in E$.
+Agents interact within a self-organized network through bidirectional communication to exchange information and update their states to align with system goals.
+Table III summarizes the notations used throughout the paper to formalize the network model and optimization framework.
+TABLE III: Nomenclature utilized in this research
 Notation Description
@@
 […]
-Agents balance their individual preferences with system-wide goals, which often involve conflicting criteria.
-Each agent has an individual preference for its plans, calculated by the discomfort cost ($D$), such that $D$=${}_{a,i}=f_{D}(p_{\mathsf{a,i}})$, where $f_{D}$ is the agent’s preference associated with a particular plan.
-Each agent $a$ measures the costs $D$ for each possible plan $p_{\mathsf{a,i}}$$\in$ $P_{\mathsf{a}}$ and the plan with the minimum cost is selected.
+Agents aim to balance their individual preferences with system-wide goals, which often involve conflicting criteria.
+Each agent $a$ has an individual preference for its plans, quantified by the discomfort cost ($D$), such that $D$=${}_{a,i}=f_{D}(p_{\mathsf{a,i}})$, where $f_{D}$ measures how undesirable plan $i$ is for agent $a$ based on the agent’s preferences; lower costs indicate more preferred plans.
+Each agent $a$ evaluates the costs $D$ for each possible plan $p_{\mathsf{a,i}}$$\in$ $P_{\mathsf{a}}$ and the plan with the minimum cost is selected.
 While agents aim to minimize their own discomfort, they may also consider system-wide metrics such as the inefficiency cost ($I$).
@@
 It represents the system-wide performance inefficiency that agents aim to minimize through coordinated decision-making: $I=f_{I}\left(\sum_{a=1}^{\mathsf{n}}(s_{\mathsf{a}})\right)$.
-Agents optimize their decisions by minimizing a weighted combination of these costs, as illustrated in Equation 1.
-The behavior of agent $a$ is modeled by the corresponding weights $\beta_{a},\text{ and }\alpha_{a}$, prioritizing efficiency and comfort, respectively.
+Each agent selects a plan that minimizes a weighted combination of individual discomfort and system-wide inefficiency, as shown in Equation 1.
+$s_{\mathsf{a}}$ $=\arg\min_{i=1}^{k}\left(\alpha_{a}\cdot I_{a,i}+\beta_{a}\cdot D_{a,i}\right)$
+$=\arg\min_{i=1}^{k}\left[\alpha_{a}\cdot f_{I}\left(s_{1}+s_{2}+\cdots+s_{\mathsf{n}}\right)\right.$
++ $\beta_{a}\cdot f_{D}\left(D_{1,s},D_{2,s},\ldots,D_{\mathsf{n},s}\right)\right]\text{,}$ (1)
+$\text{where}\quad\alpha_{a}+\beta_{a}$ $=1\quad\text{\&}\quad\alpha_{a},\beta_{a}\in[0,1]$
+The behavior of agent $a$ is modeled by the corresponding weights $\alpha_{a}$ and $\beta_{a}$, which represent agent’s priorities between minimizing system-wide inefficiency and personal discomfort, respectively.
 A higher weight indicates a higher preference for minimizing the corresponding objective.
 On the other hand, a weight of 0 means that the corresponding objective is not considered.
-For example, When $\beta_{a}=0,\text{ and }\alpha_{a}=1$, the agent prioritizes system-wide efficiency, acting as altruistic.
-Conversely, when $\beta_{a}=1,\text{ and }\alpha_{a}=0$, the agent prioritizes its own interest, behaving as selfish.
-$s_{\mathsf{a}}$ $=\arg\min_{i=1}^{k}\left(\alpha_{a}\cdot I_{a,i}+\beta_{a}\cdot D_{a,i}\right)$
-$=\arg\min_{i=1}^{k}\alpha_{a}\cdot f_{I}\left(s_{1},s_{2},\ldots,s_{\mathsf{n}}\right)$
-$\quad+\beta_{a}\cdot f_{D}\left(D_{1,s},D_{2,s},\ldots,D_{\mathsf{n},s}\right)$ (1)
-$\text{where}\quad\alpha_{a}+\beta_{a}$ $=1\quad\text{\&}\quad\alpha_{a},\beta_{a}\in[0,1]$
-The decision-making process in distributed optimization is influenced not only by each agent’s individual objectives but also by the collective impact of their choices on system-wide goals.
-This interdependence introduces significant complexity, particularly when cost functions are non-linear, as agents’ choices depend on one another.
-Minimizing inefficiency and discomfort costs under such conditions is an NP-hard problem [49].
-These challenges are further exacerbated by the lack of complete information about other agents’ selections and the need for coordination among agents’ choices.
-Addressing these issues requires a decentralized optimization framework that enables agents to iteratively refine their plan selections based on aggregated contributions and prior iterations (see Section IV-A).
+For instance, an agent with $\alpha_{a}=1$ and $\beta_{a}=0$ behaves altruistically, prioritizing global goals.
+Conversely, $\alpha_{a}=0$ and $\beta_{a}=1$ define a selfish agent focusing solely on its individual preference.
 III-D Adversarial Distributed Optimization Model
-To study the resilience of distributed optimization in an adversarial environment, we partition the set of agents, $A$, into two subsets: adversary agents ($A_{\mathsf{d}}$) and legitimate agents ($A_{\mathsf{l}}$).
-Adversary agents, $A_{\mathsf{d}}\subseteq A$, prioritize their individual interests over collective system goals by adapting their behavior to maximize personal benefits.
-This behavior disregards system-wide efficiency and leads to suboptimal outcomes for the overall system.
-Legitimate agents, $A_{\mathsf{l}}\subseteq A$, align their actions with system-wide objectives to optimize overall performance.
-The total set of agents is defined as $A=A_{\mathsf{d}}\cup A_{\mathsf{l}}$.
+We propose an adversarial model applicable across a range of combinatorial optimization scenarios.
+In this model, the agent population $A$ is partitioned into two disjoint subsets: legitimate agents $A_{\mathsf{l}}$ and adversary agents $A_{\mathsf{d}}$, such that $A=A_{\mathsf{l}}\cup A_{\mathsf{d}}$.
+While legitimate agents, $A_{\mathsf{l}}\subseteq A$, align their actions with system-wide objectives to optimize overall performance, adversary agents, $A_{\mathsf{d}}\subseteq A$, prioritize their individual interests over collective system goals by adapting their behavior to maximize personal benefits.
+For instance, in a bike-sharing system [53], optimization ensures a balanced distribution of bikes across stations to meet user demand.
+Legitimate users may select pick-up and drop-off stations while considering system-wide efficiency, maintaining network equilibrium.
+In contrast, adversary users prioritize their own convenience, selecting stations solely based on personal preference, leading to imbalances such as empty or overloaded stations, and ultimately degrading overall efficiency and user satisfaction.
+Adversarial behavior is modeled by adjusting the weight $\beta_{a}$ in the agent’s decision function (Equation 1).
+Legitimate agents are assigned $\beta_{l}=0$, fully aligning with system goals, while adversary agents are assigned $\beta_{d}>0$, increasing emphasis on personal discomfort minimization at the expense of system-wide efficiency.
+This parameterization allows adversarial behavior to be modeled in a continuous space, from fully altruistic to fully selfish.
+By varying the distribution and severity of adversarial weights across the agent population, our model enables systematic analysis of resilience, vulnerability, and collapse in distributed optimization.
+This includes assessing how the impact of adversarial behavior is linked to the network positioning of the adversary agents and legitimate agents.
 Adversarial behavior amplifies the inefficiency cost $I$, reflecting the trade-off between minimizing individual discomfort ($D$) and optimizing overall system efficiency ($I$).
-While $D$ focuses on individual preferences, $I$ addresses the system-wide inefficiency caused by deviations from optimal resource allocation, underscoring the conflict between personal and collective goals.
-For instance, in a bike-sharing system, optimization ensures a balanced distribution of bikes across stations to meet user demand.
-Legitimate users select pick-up and drop-off stations with consideration for system-wide efficiency, aiming to maintain network equilibrium.
-Conversely, adversary users prioritize their convenience, choosing stations that minimize their discomfort without regard for the network state.
-This behavior creates imbalances, such as empty or overstocked stations, reducing overall efficiency and user satisfaction.
+While $D$ focuses on individual preferences, $I$ addresses the system-wide inefficiency caused by deviations from optimal resource allocation, underscoring the conflict between individual and collective optimization goals.
 IV Experimental Methodology
+This section illustrates the distributed optimization method employed as a case study, the experimental setup, the application scenarios, the measured variables and the evaluation metrics.
 This paper proposes a novel adversarial distributed optimization model to study the performance of multi-agent discrete-choice combinatorial optimization under distributed adversary attacks.
-IV-A Decentralized Optimization Framework
-The proposed model is implemented within the Iterative Economic Planning and Optimized Selections (I-EPOS) [^2] Available at https://github.com/epournaras/EPOS, a collective learning framework.
-I-EPOS employs a self-organized, multi-level hierarchical network structure that ensures efficient communication and coordinated decision-making while reducing communication overhead [51].
-I-EPOS enables the agents to iteratively adjust their plan selections based on the aggregated plans of the agents beneath and previous iteration choices.
-The proposed model operates iteratively, enabling agents to coordinate their choices in collective decision-making.
+IV-A Distributed Optimization Method
+The adversarial distributed optimization model is implemented within the Iterative Economic Planning and Optimized Selections (I-EPOS) framework.
+I-EPOS is a discrete-choice distributed combinatorial optimization algorithm for large-scale multi-agent networks [53, 54].
+It employs a self-organized, multi-level hierarchical structure to enable efficient communication, coordination, and scalability while minimizing communication overhead [57].
+I-EPOS enables the agents to iteratively coordinate their choices in collective decision-making.
 Each iteration consists of two distinct phases: a bottom-up phase and a top-down phase.
-During the bottom-up phase, agents select plans based on the aggregated choices of agents in the subtree beneath, as well as the selections made by all agents in the previous iteration.
+During the bottom-up phase, agents select plans based on the aggregated choices of agents in the branch beneath, as well as the selections made by all agents in the previous iteration.
 Conversely, the top-down phase addresses incomplete knowledge from higher branches in the hierarchy, enabling agents to revert to previous selections if no cost reduction is achieved.
 This process continues until a predefined iteration limit is reached or no further improvement in the optimization objective occurs [53, 54].
-The hierarchical network structure is organized as an acyclic graph, where each parent agent aggregates responses from its children without redundancy, explicitly avoiding double counting of contributions from lower nodes in the hierarchy.
-This design ensures efficient coordination and alignment of individual decisions with system-wide objectives while maintaining overall optimization [49, 51, 52].
-I-EPOS is well-suited for adversarial scenarios in Smart City applications due to its scalability, adaptability to diverse agent behaviors, and ability to handle adversarial conditions [50, 53].
+This iterative coordination mechanism addresses the inherent complexity of multi-agent optimization, particularly under non-linear cost functions and incomplete knowledge of other agents’ choices.
+These conditions make the optimization problem NP-hard [53], requiring distributed coordination methods that allow agents to refine their decisions based on both local preferences and system-wide impact.
+The hierarchical network is structured as an acyclic graph, where each parent agent aggregates responses from its children by avoiding double counting.
+This design ensures efficient coordination when optimizing individual decisions and system-wide objectives [53, 57, 55].
+I-EPOS is well-suited for adversarial scenarios in large-scale distributed systems due to its scalability, adaptability to diverse agent behaviors, and potential mitigate adversarial conditions [54, 58].
+To apply the proposed adversarial model, the I-EPOS framework was extended to support heterogeneous agent behaviors.
+The original implementation assumed uniform agent preferences across the population.
+We enhanced the framework to allow agents to configure individual decision-making weights.
+This enhancement enables modeling adversarial agents with varying levels of behavioral severity and is made available as an open-source artifact to support reproducibility and future research [^4] Available at https://github.com/epournaras/EPOS..
 IV-B Experimental Setup
-Experiments are conducted using multiple HPC servers with varying configurations that support large-scale simulations and ensure computational efficiency.
+Experiments are conducted using multiple HPC servers with varying configurations that support large-scale experimentation and ensure computational efficiency.
 These include high-memory nodes (up to 768 GB) and multi-core processors (up to 40 cores per node).
@@
 IV-C Application Scenarios
-TABLE III: Description of the datasets and experimental setup in this research
+Adversarial distributed optimization is studied in three application scenarios based on real-world data and a synethic dataset.
+Table IV provides an overview of the datasets used in this research, including the agent populations, the number and size of plans per agent, agent representation and the interpretation of discomfort and inefficiency costs within each application domain.
+These costs are defined explicitly through the optimization objectives and reflect domain-specific constraints.
+Further mathematical definitions of cost functions are provided in Appendix A: Cost Function Definitions.
+TABLE IV: Description of the datasets and experimental setup in this research
 Dataset Name No.
 Agents No.
-Plans Plan Size Agents Discomfort Cost Inefficiency Cost Total Experiments Energy 1000 10 144 Households Agent preference based on time shift Energy demand 30,018,000 Data-collective 72 3 64 participants Privacy loss Data collective inefficiency (quality of collected data) 65,433,600 (2 target signals) Voting 266 31 5 Voters Compromise Polarization 97,551,840 (120 target signals)
+Plans Plan Size Agents Discomfort Cost Inefficiency Cost Total Experiments Energy 1000 10 144 Households Time shift from intrinsic preference Variance of energy demand 3,118,560 Privacy 72 3 64 Smart phone users Privacy loss Mismatch between shared and desired data 498,780 (2 target signals) Voting 266 31 5 Voters Compromise distance from intrinsic voting preferences Polarization 103,456,800 (120 target signals) Gaussian (synthetic) 10–100 2–10 2 Simulated agents Ranking distance Variance 4,125,000
 IV-C1 Energy-demand dataset
@@
 The privacy dataset originates from a living-lab experiment at the Decision Science Laboratory [^7] https://www.descil.ethz.ch of ETH Zurich, involving 72 participants evaluating 64 data-sharing scenarios that involve 4 sensor types, data collectors, and contexts [58].
-The data-sharing choices of each participant determine three data-sharing plans, representing their intrinsic motivation to share and two rewarded scenarios.
+The data-sharing choices of each participant in the experiment determine three data-sharing plans, representing their intrinsic motivation to share and two rewarded scenarios.
 The plans are assessed using privacy valuation schemes assigning normalized costs ranging from 0 to 1, where lower costs indicate less privacy compromise.
-The dataset facilitates testing under high and low privacy-preservation target signals, reflecting intrinsic preferences for data-sharing behaviors.
-The inefficiency cost is calculated based on the reduction in the quality of the aggregated data due to adversarial participants prioritizing minimal data sharing over collective goals.
+The dataset facilitates testing under high and low privacy-preservation target signals.
+The inefficiency cost is calculated by the residual sum of squares between the shared and the desired data that measures their mismatch and is an indicator of quality of service supported by the collected data.
 Adversarial participants disrupt coordination by focusing solely on minimizing their data sharing, under-mining the quality of service of data collectors.
@@
 Each voter has 31 alternative voting plans, representing ranked preferences among the five candidates.
-The optimization focuses on minimizing polarization in decision outcomes, which refers to the extent of divergence in voters’ preferences from a collectively agreed compromise.
-Inefficiency cost is measured as the cost of polarized decision outcomes.
-To ensure fair representation, 120 target signals are generated from all combinations of values 0, 0.25, 0.5, 0.75, and 1.
-Adversarial behavior occurs when voters prioritize minimizing their personal compromise cost, disregarding collective consensus, which amplifies polarization and disrupts the system ability to achieve shared, balanced outcomes.
-Table III provides an overview of each dataset used in this research, including the number of agents, the number and size of plans per agent, agent representation, the corresponding discomfort and inefficiency cost functions, and the total number of experiments conducted.
+The optimization focuses on minimizing polarization in the voting outcomes, which refers to reaching the same voting outcome but with compromises that reduce polarization.
+Polarization here is the inefficiency cost and it refers to the mismatch from a liner ranking of the alternatives in the voting outcomes, although other polarization models could be studied as well [60].
+The rationale of linearity is to deviate from concentrating the voters’ preferences to two opposing poles.
+To control for the same voting outcome, 120 target signals are generated from all combinations of values 0, 0.25, 0.5, 0.75, and 1, representing the linear ranking of alternatives.
+Adversarial behavior occurs when voters prioritize their intrinsic preferences, i.e., no compromises to reduce polarization.
+IV-C4 Synthetic Gaussian dataset
+The synthetic dataset is constructed to evaluate system scalability under controlled, domain-agnostic conditions.
+It includes 100 agents with 10 generated plans, where each plan is a 100-dimensional vector sampled from a Gaussian distribution $\mathcal{N}(0,1)$.
+Plans are sorted by their index, with lower indices arbitrary indicating higher agent preference.
+Discomfort cost is defined as the rank of the selected plan—i.e., a higher index reflects greater deviation from the most preferred option.
+Inefficiency cost is measured as the variance of the aggregated global response, capturing system-wide imbalance.
+This synthetic setup allows systematic analysis of adversarial effects across varying agent populations, number/size of plans, and attack configurations.
 IV-D Varying Dimensions and Performed Experiments
-The experiments systematically evaluate the resilience and vulnerability of the system across the following dimensions:
+The following dimensions are studied in the performed experiments:
   • Scale of adversaries ($|A_{\mathsf{d}}|$): Incrementally increase the number of adversary agents $A_{\mathsf{d}}$ from 1 to $\mathsf{n}$ across all datasets to analyze performance under varying adversary densities; i.e., $A_{\mathsf{d}}=\{a\mid a\in A\}$.
-  • Adversarial severity ($\beta$): The adversarial preference to minimize discomfort cost ($\beta$) is varied across 30 levels, with $\beta$ incrementing from 0 to 1 such as $\beta=\frac{b}{30}$ for $b=\{1,2,3,...,30\}$.
-This variation captures the impact of varying levels of adversarial intent, where agents prioritize their personal objectives at the expense of system-wide goals, on overall system performance.
+  • Adversarial severity ($\beta$): The adversarial preference to minimize discomfort cost ($\beta$) is varied across 30 levels, with $\beta$ incrementing from 0 to 1 such that $\beta=\frac{b}{30}$ for $b=\{1,2,3,...,30\}$.
   • Adversary position: The influence of adversary positions within the hierarchical network is analyzed using two approaches: layer-wise and cumulative structural analysis.
 A binary tree structure is employed, with each hierarchical layer containing approximately $\log_{2}|A|$ agents, where $|A|$ is the total number of agents.
-The layer-wise analysis evaluates inefficiency costs across individual hierarchical layers under varying adversary scales (25%, 50%, 75%, 100%) and behavior levels ($\beta$ = 0.1, 0.4, 0.7, 1.0).
-The cumulative analysis examines the aggregated impact of adversary agents propagating through the network in top-down (root-to-leaf) and bottom-up (leaf-to-root) configurations.
-[…]
+The structural analysis evaluates inefficiency costs under varying adversary scales (25%, 50%, 75%, 100%) at each layer of the hierarchy.
+The cumulative analysis examines the aggregated impact of adversary agents positioned incrementally in top-down (root-to-leaf) and bottom-up (leaf-to-root) configurations.
+For each dataset, experiments are conducted across 30 adversarial severity levels ($\beta$) with 100 simulation runs per configuration.
+Layer-wise structural analysis is performed at four adversarial proportions $p\in\{25\%,50\%,75\%,100\%\}$ within each layer of the hierarchical topology, where the number of layers is defined as $\lceil\log_{2}|A|\rceil$.
+For each layer $L$, the number of adversarial agents is calculated as $k_{p}=\max\left(1,\ \left\lceil\frac{p}{100}\cdot|A_{L}|\right\rceil\right)$, where $|A_{L}|$ denotes the number of agents in layer $L$.
+Adversarial configurations are sampled up to 100 combinations per setting to ensure computational feasibility [^9] The equation assumes up to 100 combinations per layer-percentage, but the binary hierarchy often yields fewer due to limited agents per layer..
+In addition, two cumulative structural attack scenarios are simulated, top-down (root-to-leaf) and bottom-up (leaf-to-root), introducing $2\times|A|$ further experiments per dataset.
+The total number of experiments per dataset is calculated as:
+Total Experiments=(30×number of $\text{Total Experiments}=(30\times\text{number of signals})\times\left[(100\times|A|)+\sum_{L}\sum_{p}\min\left(100,\ \binom{|A_{L}|}{k_{p}}\right)+(2\times|A|)\right]$
+For the synthetic Gaussian dataset, experiments vary both agent populations and the number of plans per agent (from 2 to 10).
+The total number of experiments is computed as: $\text{Total Experiments}_{\text{Gaussian}}=\sum_{i=1}^{10}\left(10i\times 30\times 5\times 50\right),$ where $i$ is the number of agents, 30 is the number of severity levels, 5 is the number of plans, and 50 is the number of random structural permutations (i.e., reordering of agents in the tree).
 IV-E Evaluation Metrics
-Optimization objectives: Three metrics evaluate system performance: inefficiency cost $I$, discomfort cost $D$, and compromise cost of legitimate agents.
-The inefficiency cost measures the overall cost incurred by the system.
-The discomfort cost evaluates agents’ dissatisfaction with non-preferred choices.
-The compromise cost quantifies the extent to which legitimate agents adjust their individual preferences to align with collective goals under adversarial influence.
-This metric is calculated as the deviation between the selected plans of legitimate agents and their most preferred plans.
-These metrics evaluate the trade-offs between individual agent goals and system-wide objectives, particularly under varying adversarial conditions.
-Pareto optimality: In multi-objective optimization, the Pareto front represents the set of non-dominated solutions where no objective can be improved without compromising another.
-The knee point on the Pareto front indicates the solution that achieves the best trade-off between competing objectives, making it a key decision-making criterion [55, 56].
-This study identifies the knee point using the Minimum Manhattan Distance (MMD) method, which measures the distance between each solution on the Pareto front and an ideal reference point where both objectives are optimized.
-The solution with the smallest distance to the ideal point is selected as the knee point.
-This approach ensures a balanced evaluation of discomfort cost ($D$) and inefficiency cost ($I$) while maintaining alignment with established methodologies in the literature [56].
-Resilience, vulnerability, and collapse framework: To define thresholds for resilience, vulnerability, and collapse in distributed optimization, we apply the multi-Otsu thresholding method.
-The Otsu method is an efficient technique for dividing data into distinct regions by minimizing intra-class variance [57, 58].
-In this study, the method segments data into three regions based on inefficiency levels: resilience (low inefficiency), vulnerability (moderate inefficiency), and collapse (high inefficiency).
-This framework provides a robust method for segmenting inefficiency data into resilience, vulnerability, and collapse states, enabling a clear assessment of system performance transitions under adversarial conditions.
-V Results Analysis and Discussion
-This section presents the results of extensive experiments evaluating the impact of adversarial agents on multi-objective distributed optimization across the Energy, Voting, and Privacy datasets.
+Optimization objectives: System performance is assessed using three metrics: inefficiency cost $I$, discomfort cost $D$, and the compromise cost of legitimate agents.
+Inefficiency cost captures the deviation from optimal system performance.
+Discomfort cost reflects individual agents’ dissatisfaction, i.e.
+to what extent a plan is not the most preferred one.
+The compromise cost quantifies the increase in discomfort experienced by legitimate agents due to adversarial influence.
+It is calculated as the difference in discomfort between scenarios with and without adversaries, highlighting the collective burden legitimate agents bear to maintain system performance.
+Pareto optimality: In multi-objective optimization, the Pareto front defines solutions where no objective can improve without compromising another objective.
+The knee point on this front identifies the most balanced trade-off between competing objectives.
+This study uses the Minimum Manhattan Distance (MMD) method to locate the knee point, measuring the distance from each Pareto solution to an ideal reference point where both objectives are optimized.
+The solution with the smallest distance is selected.
+This approach ensures a balanced trade-off between discomfort ($D$) and inefficiency ($I$) in line with established approaches [61, 62].
+Resilience, vulnerability, and collapse framework: To classify system states, the multi-Otsu thresholding method is applied to segment inefficiency and discomfort values into three distinct regions: resilience (low inefficiency), vulnerability (moderate inefficiency), and collapse (high inefficiency).
+This technique minimizes intra-class variance, offering a robust framework to detect transitions in system performance under adversarial conditions [63, 64].
 (a) Inefficiency costs
 (b) Discomfort costs
-(c) Compromised discomfort costs for non-adversary agents
-Figure 1: Multi-objective optimization performance across different numbers of adversary agents and various levels of adversaries within the Energy, Voting, and Privacy datasets.
-R, V, and C indicate regions of Resilience, Vulnerability, and Collapse.
-V-A Resilience Analysis
-Figure 1 illustrates inefficiency costs, discomfort costs, and compromised discomfort costs across varying scale of adversaries and behavior levels.
-These metrics collectively capture system-wide inefficiency, individual agent dissatisfaction, and the trade-offs faced by legitimate agents under adversarial conditions.
-Energy Dataset
-[…]
-Voting Dataset
-The Voting dataset demonstrates significant stability, maintaining resilience across more than 90% of the operational space.
-Inefficiency remains low for $\beta<0.9$ and scale of adversaries below 50%.
-Vulnerability emerges between 50% and 70% adversary agents, with inefficiency increasing by 111%.
-Collapse occurs at $\beta\geq 0.96$ and proportions exceeding 70%.
-Discomfort costs decline gradually in resilient regions and are entirely eliminated under collapse conditions, highlighting a clear trade-off between individual comfort and system-wide efficiency.
-Compromised discomfort costs show resilience for adversary agents below 20%, remaining low even under high levels of adversarial behavior.
-However, collapse is observed for proportions exceeding 50% and adversarial levels $\beta>0.7$, where compromised discomfort costs rise smoothly.
-This trend suggests that legitimate agents in the Voting dataset show lower resilience compared to the Energy dataset.
-Privacy Dataset
-The Privacy dataset reveals lower resilience thresholds compared to the Energy and Voting datasets, especially for high privacy-preserving signal.
-For $\beta<0.2$ and adversary agents below 20%, inefficiency costs remain low.
-However, vulnerability emerges as proportions increase to 50%, and collapse occurs at $\beta\geq 0.3$ for high data-sharing levels and $\beta\geq 0.5$ for low data-sharing levels when adversary agents exceed 50%.
-Discomfort costs follow similar trends, with resilience observed below 30% adversary agents for high privacy-preserving signal and 20% for low privacy-preserving signal.
-Collapse occurs when $\beta>0.3$ and proportions exceed 60%.
-Compromised discomfort costs rise steadily, peaking at values of 9 and 10 for high and low privacy-preserving signal, respectively, at 90% adversary agents.
-Unlike the Energy and Voting datasets, where $\beta$ strongly influences compromised discomfort costs, the Privacy dataset shows that scale of adversaries serves as the dominant driver of these costs beyond $\beta>0.3$.
-This highlights the Privacy dataset heightened sensitivity to adversarial density, requiring more targeted interventions to preserve legitimate agent objectives.
+(c) Compromised discomfort costs for legitimate agents
+Figure 1: Inefficiency, discomfort, and compromised discomfort over adversary scales and severity in the energy, voting, and privacy datasets, , including Pareto knee points and resilience (R), vulnerability (V), and collapse (C) thresholds.
+V Results Analysis and Discussion
+This section presents the results of extensive experiments evaluating the impact of adversarial agents on multi-objective distributed optimization across the real‐world (energy, voting, and privacy) and synthetic datasets.
 (a) Pareto front and knee points across adversarial severity.
@@
 Figure 2: The Pareto Optimality of The energy voting, and privacy datasets
+V-A Resilience Analysis
+Figure 1 presents inefficiency, discomfort, and compromised discomfort costs across varying adversary scales and severity levels ($\beta$).
+These metrics capture degradation in system performance, agent satisfaction, and the impact on legitimate agents.
+The graphs include resilience, vulnerability, and collapse thresholds, with overlaid Pareto knee points.
+For the inefficiency and compromised‐discomfort plots, knees mark the best trade‐off between those two metrics; on the discomfort heatmaps, the knee indicates the minimal total discomfort for a given inefficiency.
+Inefficiency cost (Figure 1(a)) remains low in the resilience zone, especially when adversary ratios are below 30–50% and $\beta<0.8$.
+As adversary scale and severity increase, systems shift from Resilience to Vulnerability and Collapse, with thresholds varying across datasets.
+Energy and voting maintain  90% resilience and only 3–5% collapse, while privacy configurations show earlier collapse near 20%.
+In energy, inefficiency peaks at $\beta=1$ near 4000 cost  [^10] Values for $\beta=1$ in the energy dataset are excluded from visualizations to avoid heatmap saturation due to extremely high inefficiency values, when adversaries exceed 85%.
+In the voting dataset, inefficiency increases by 111% in the vulnerable region, with collapse triggered beyond 70% adversaries at $\beta\geq 0.96$.
+Privacy collapses earlier: $\beta\geq 0.3$ at 50% adversaries in the high privacy-preserving signal with 50% adversaries, and $\beta\geq 0.5$ in the low privacy-preserving signal.
+Discomfort cost (Figure 1(b)) shows a consistent decline as adversarial intensity increases.
+In resilient regions, typically below 30–50% adversary presence and $\beta<0.7$, discomfort initially remains high but declines gradually, then sharply in collapse regions, as adversary scale and severity increase.
+In the energy dataset, discomfort remains high only below 30% adversaries and drops by 8% before collapsing to near-zero.
+The voting dataset follows a similar pattern, with discomfort gradually decreasing and fully eliminated in collapse.
+The privacy datasets exhibit earlier collapse ($\beta>0.3$, $>$60% adversaries), resulting in faster discomfort decline.
+Compromised discomfort cost (Figure 1(c)) increases with adversarial presence.
+In the energy dataset, it remains low for $\beta<0.8$ even when all agents are adversarial, but rises sharply when the adversary scale exceeds 40% at $\beta>0.9$.
+In voting, the cost remain low under mild attacks ($\beta<0.2$, adversary scales $<20\%$) and increase gradually beyond 50% at $\beta>0.7$.
+Privacy datasets show a steeper rise, peaking at 9 (high privacy) and 10 (low privacy) when 90% of agents are adversarial.
+Resilience disappears at high adversary densities ($>80\%$), even under low severity, and collapse emerges at low $\beta$ when agent compromise is widespread.
+While energy and voting are more sensitive to $\beta$, privacy is primarily affected by adversary scale.
+Interestingly, a noticeable resilience lag exists between the system-level inefficiency and the agent-level discomfort and compromise costs.
+In all datasets, a substantial portion of configurations classified as ”Vulnerable” or even ”Collapsed” by inefficiency remain in the ”Resilient” state when evaluated by discomfort metrics.
+For instance, in the energy dataset, while only 6% of configurations are in vulnerability based on inefficiency, 59% fall under vulnerability based on discomfort—indicating that discomfort degrades far earlier.
+Similarly, in the voting dataset, 21% of configurations are classified as vulnerable by inefficiency, compared to 70% by discomfort.
+This pattern reveals a lag of up to 40–60% in the transition from individual to system-level degradation, with discomfort and compromise metrics acting as early-warning indicators long before global inefficiency surfaces.
+This insight underlines the importance of incorporating agent-centric metrics for proactive resilience monitoring.
 (a) Energy dataset.
@@
 V-B Pareto Optimality Analysis
-Pareto optimality analysis identifies Pareto fronts knee points that represent optimal system performance under varying adversarial conditions.
-Figure 2 highlights system resilience to adversarial severity and proportions across the Energy, Voting, and Privacy datasets.
-Figure 2(a) focuses on the impact of adversarial severity ($\beta$) on tolerated scale of adversaries, while Figure 2(b) explores how varying adversary densities affect tolerated behavior levels.
-[…]
-The Voting dataset demonstrates robust stability.
-Figure 2(a) shows consistent knee points tolerating 50–60% adversary agents across a broad range of behavior levels ($0.1\leq\beta\leq 0.7$).
-Even under high adversarial severity ($0.86\leq\beta\leq 0.9$), as illustrated in Figure 2(b), the system exhibits exceptional resilience, tolerating densities from 10% to 100%.
-This highlights the dataset high tolerance to adversarial influence.
-The Privacy dataset exhibits slightly difference dynamics between high and low privacy-preserving signals.
-In Figure 2(a), tolerated scale of adversaries for high signal configurations are approximately 70% at $\beta<0.5$, but drop to 50% at higher intensities.
-The low signal Privacy scenario displays smoother declines, with tolerated densities averaging 68% Adversary agents across behavior levels.
-Figure 2(b) shows that high signal configurations begin at knee points of $\beta=0.03$ for 10% adversary agents, rising gradually to $\beta=0.2$ at 50–90%.
-Low privacy-preserving signal exhibits a smoother gradient, peaking at $\beta=0.36$ at 20% Adversary agents and declining to $\beta=0.13$ at full adversarial saturation.
+Pareto optimality analysis identifies fronts and knee points reflecting optimal trade-offs under varying adversarial conditions.
+Figure 2 shows how system inefficiency relates to discomfort experienced by legitimate agents across different severity levels and adversary scales.
+Figure 2(a) focuses on the impact of adversarial severity ($\beta$) on the tolerated scale of adversaries.
+The voting dataset maintains stable knee points across 50–60% adversary ratios over a broad severity range ($0.1\leq\beta\leq 0.7$), with consistent fronts even at high $\beta$.
+The privacy dataset shows distinct patterns: the high privacy-preserving signal tolerates 70% adversaries at $\beta<0.5$, decreasing to 50% at higher severities; the low privacy-preserving signal exhibits smoother transitions with average tolerance near 68%.
+Figure 2(b) evaluates the impact of adversarial scale on tolerated severity.
+For visual clarity, only 20 adversary scales are shown per dataset.
+In energy, knee points are stable at $0.30.03<\beta<$ 0.3 for up to 90% adversaries, increase to $\beta=0.9$ at lower scales ($<40\%$), and decline to $\beta<0.1$ at full scale.
+Voting shows higher tolerance, with knee points extending $\beta=0.8$ even under full adversaries.
+Privacy dataset shows greater variability: the high privacy-preserving configuration reaches $\beta=0.2$ for 50–90% adversaries, while the low privacy-preserving case peaks at $\beta=0.36$ for 20% adversaries and declines to $\beta=0.13$ at full scale.
+Detailed Pareto plots are in Appendix Appendix B: Pareto Optimality Visualizations.
+Figure 4: Inefficiency cost across hierarchical structure layers under various adversarial configurations in energy, voting and privacy datasets.
 V-C Structure Analysis
-This section analyzes how hierarchical structures influence inefficiency costs in Energy, Voting, and Privacy datasets.
-This analysis is divided into two approaches: individual layer-wise and cumulative structure analysis.
+This section analyzes how hierarchical structures influence inefficiency costs across datasets using two approaches: layer-wise and cumulative structural analysis.
 V-C1 Layer-wise Structural Analysis
-Figure 3 highlights inefficiency costs across hierarchical layers for varying scale of adversaries at 25%, 50%, 75%, and 100% and behavior levels ($\beta$) at 0.1, 0.4, 0.7, and 1.0.
-In all datasets, deeper layers and layers with higher agent proportions exhibit greater inefficiency escalation under intensified adversarial influence.
-In the Energy dataset (Figure 3(a)), with 1,000 agents distributed across 10 layers, inefficiency costs remain stable up to Layer 5 at moderate adversarial severity ($\beta=0.1$, 0.4, and 0.7).
-Beyond Layer 6, costs fluctuate and peak at Layers 7 and 8.
-Notably, while Layer 9 contains the highest number of agents (256), inefficiency only peaks at $\beta=1.0$ and 100% Adversary agents, highlighting the compounded vulnerability of deeper layers under extreme adversarial influence.
-This emphasizes the increasing impact of both adversary density and adversarial severity on deeper layers.
-The Voting dataset (Figure 3(b)), with 266 agents distributed across 9 layers, shows minimal inefficiency in the upper layers (1–5) across all adversarial configurations.
-However, inefficiency rises sharply beyond Layer 5, with Layer 7 consistently incurring the highest costs, especially at higher adversarial severity ($\beta\geq 0.7$) and scale of adversaries of 75% and 100%.
-At $\beta=1.0$, inefficiency increases steadily from the root (Layer 1) to the leaves (Layer 9), highlighting the compounding impact of adversarial behavior as it propagates deeper into the hierarchy.
-The Privacy dataset (Figures 3(c) and 3(d)), with 72 agents distributed across 7 layers, exhibits almost similar trends for high and low privacy-preserving signal.
-At low adversarial behavior ($\beta=0.1$), inefficiency costs remain stable across all layers, demonstrating resilience to mild adversarial influence.
-At higher levels of adversarial behavior ($\beta\geq 0.7$), inefficiency peaks at Layer 6, which has the highest agent proportion (32 agents).
-Interestingly, Layer 7, with fewer agents, incurs lower costs even under severe adversarial conditions.
-Minor deviations, such as Layer 5 occasionally surpassing Layer 6 in the low signal scenario, highlight the nuanced relationship between layer depth, agent density, and adversarial behavior.
-These results suggest that inefficiency costs in hierarchical systems are highly sensitive to layer depth, scale of adversaries, and the level of adversarial behavior.
-Deeper layers and those with higher agent proportion emerge as critical vulnerabilities under intensified adversarial configurations.
-Reinforcing deeper layers and redistributing agent densities could mitigate inefficiency spikes, especially in applications with high adversarial risks.
-(a) Top-down cumulative
-(b) Bottom-up cumulative
-Figure 4: Inefficiency cost across hierarchical structure layers under various adversarial configurations in Energy, voting and Privacy datasets.
+Figure 3 presents the inefficiency costs across hierarchical layers under adversary scales of 25%, 50%, 75%, and 100%.
+For layers with few agents, intermediate adversary ratios are approximated by averaging to the closest feasible configurations.
+In the energy dataset (Figure 3(a)), with 1,000 agents over 10 layers, inefficiency remains low (below 0.16) across moderate adversarial scales and gradually increases with severity ($\beta<1.0$).
+The minimum inefficiency occurs at layer 2 under 100% adversaries and $\beta=0.56$, while inefficiency peaks at layer 8 under $\beta=0.96$, marking a 144% increase from the minimum.
+Profiles remain smooth under 25% and 50% adversary scales but fluctuate more at 75% and 100%.
+Under extreme severity ($\beta=1$), inefficiency surges sharply, reaching over 723 at full adversarial saturation.
+With this severity, the root layer shows moderate inefficiency, exceeding layers 2–7 at 75% and matching layers 2 and 3 at 100%, followed by a steady increase down the hierarchy.
+Despite layer 10 hosting the largest agent population, its inefficiency values are slightly lower than those of layer 9.
+In the voting dataset (Figure 3(b)), with 266 agents over 9 layers, inefficiency remains stable across all layers under low adversarial severity ($\beta\leq 0.3$).
+As severity increases, costs escalate, particularly at the root (Layer 1) and Layers 2 and 3, which consistently experience higher inefficiency than deeper layers.
+The maximum inefficiency is observed at layer 8 under 100% adversaries, where agent density is highest.
+Layer 9,despite its depth, shows lower inefficiency due to a smaller agent count.
+At 25% and 50% adversary scales, inefficiency profiles remain relatively smooth; however, at 75% and 100%, cost escalations become pronounced, particularly in upper and middle layers.
+In the privacy dataset (Figures 3(c) and 3(d)), with 72 agents across 7 layers, inefficiency remains low and uniform under low severities.
+In the high privacy-preserving configuration, the root layer consistently incurs higher costs than subsequent layers across all adversary scales, with inefficiency peaking at layer 6 (the layer with the largest number of agents) under $\beta\geq 0.7$.
+Layer 7, although deeper, has lower costs due to fewer agents.
+In the low privacy-preserving configuration, similar trends are observed with occasional fluctuations between layers 5 and 6 at higher adversary scales.
+Overall, inefficiency profiles remain structurally consistent between both privacy settings, though severity levels accelerate cost increases in the high-privacy case.
 V-C2 Cumulative Structural Analysis
-Another approach assesses the cumulative impact of adversary agents on inefficiency costs as they propagate through the hierarchical network structure in two configurations: top-down (root-to-leaf) and bottom-up (leaf-to-root).
-Figure 4 presents inefficiency trends for scale of adversaries and adversarial severity (0.1, 0.4, 0.7, 1.0) across Energy, Voting, and Privacy datasets.
-The cumulative structural analysis of the Energy dataset reveals distinct differences in inefficiency trends between top-down and bottom-up propagation.
-At $\beta=1$, top-down propagation results in significantly higher initial inefficiency costs, reaching 272 with a single adversarial agent, compared to just 0.8 in bottom-up propagation.
-Inefficiency costs in the top-down configuration also rise more rapidly, indicating accelerated system deterioration.
-In contrast, bottom-up propagation exhibits a slower emergence of inefficiency, with costs remaining low up to 20% adversarial agent participation, followed by exponential growth.
-These trends emphasize the critical impact of adversarial agent placement, as costs escalate faster and peak higher when adversaries are positioned in top layers.
-At lower adversarial severity (0.1, 0.4, 0.7), inefficiency costs remain low and stable, fluctuating between 0.1 and 0.5, even as number of adversaries increases.
-This highlights the system resilience to moderate adversarial behavior.
-The Voting dataset shows distinct trends in inefficiency costs.
-At $\beta=1$, inefficiency costs start at approximately 2.8, similar to lower $\beta$ levels, but rise sharply as the proportion of adversary agents increases, peaking at 6.0 when adversary agents saturate the network (100%).
-In the top-down propagation, inefficiency costs accelerate rapidly within the top four layers, driven by early adversarial influence.
-This is followed by a clear decrease between Layers 5 and 6, corresponding to scale of adversaries ranging from 6% to 17%, indicating that inefficiencies stabilize as the system progresses deeper into the hierarchy.
-At lower adversarial severity ($\beta=0.1,0.4,0.7$), inefficiency costs remain low and stable, fluctuating slightly between 2.8 and 2.9, regardless of scale of adversaries.
-Notably, at $\beta=0.7$, inefficiency costs increase faster in the top-down propagation compared to the bottom-up propagation, underscoring the impact of adversary agents propagation direction on system performance under moderate adversarial behavior.
-The Privacy dataset reveals resilience at low adversarial severity ($\beta=0.1$), with minimal inefficiency increases (0.01–0.05) across all layers in both top-down and bottom-up propagations, even under full adversary participation.
-At higher adversarial levels ($\beta=0.7,0.4$ and $\beta=1$), inefficiency costs rise significantly across layers.
-For high privacy-preserving signal, inefficiency emerges early with just one adversary agent in top-down propagation at Layer 1 with $\beta=0.4,0.7$, and $1$, while in bottom-up propagation, inefficiency begins with low value $<62.8$.
-In both propagations Layer 2 shows heightened sensitivity, where even 2-3 adversary agents sharply increase inefficiency.
-Layer 6 experiences notable surges when the scale of adversaries exceeds 50% for both propagation directions.
-In low privacy-preserving signals, inefficiency costs rise more prominently in top-down propagation, particularly at Layer 5, with sharp increases at $\beta=0.4,0.7$, and $1$ when scale of adversaries exceeds 25%.
-In bottom-up propagation, inefficiency escalates at $\beta=0.7$ and $1$ when scale of adversaries exceeds 70%.
-These patterns emphasize structural vulnerabilities, with Layers 2, 5, and 6 emerging as critical points for inefficiency escalation under intensified adversarial influence.
-V-D Discussion
-The results of this study reveal critical insights into adversarial influence on multi-objective distributed optimization systems, uncovering opportunities for designing self-healing strategies that enhance system resilience.
-By examining the dynamics of resilience, vulnerability, and collapse across diverse datasets, the analysis offers valuable lessons for addressing challenges in adversarial environments.
-The Energy and Voting datasets exhibit strong resilience, withstanding a broad range of adversarial configurations and transitioning smoothly from vulnerability to collapse under higher adversarial severity.
-In contrast, the Privacy dataset demonstrates heightened sensitivity to adversarial scales and severity, particularly under high privacy-preservation constraints.
-This results underscore the importance of tailoring optimization strategies, particularly for privacy-preserving systems.
-Pareto analysis identifies critical resilience thresholds where systems can balance inefficiency and discomfort costs.
-The Energy dataset sustains resilience up to 90% adversarial scale, collapsing only under extreme severity, while the Voting dataset maintains consistent performance across configurations, showcasing robustness.
-The Privacy dataset requires earlier interventions for high privacy-preserving configurations but allows for more gradual adjustments under low privacy-preserving scenarios.
-These findings reinforce the importance of identifying Pareto knee points to enable dynamic trade-off prioritization and effective system adaptation under real-time adversarial conditions.
-Structural analysis reveals vulnerabilities in deeper hierarchical layers with higher number of agents, which are more prone to inefficiency propagation under intensified adversarial influence.
-Targeted strategies, such as reinforcing critical layers, redistributing agent densities, and employing localized containment mechanisms, can mitigate inefficiencies.
-Top-down configurations show rapid cascading of adversarial impact from top layers, demanding robust protections at higher levels, whereas bottom-up configurations exhibit slower propagation, suggesting inherent resilience.
-Integrating adaptive fault-monitoring mechanisms is essential for maintaining system stability.
-Dynamically detecting transitions between resilience, vulnerability, and collapse allows for strategic resource allocation.
-For instance, energy systems can employ load-balancing strategies to address inefficiency, while hierarchical vulnerabilities in distributed optimization algorithms should be addressed through structural reinforcements and proactive monitoring.
-This research provides a comprehensive framework for understanding adversarial impacts on distributed optimization systems and offers actionable insights for designing robust, self-healing mechanisms.
-By combining structural interventions, adaptive monitoring, and informed resource allocation, distributed systems can maintain efficiency and resilience even under significant adversarial configurations.
+Cumulative analysis evaluates how adversarial influence propagates through hierarchical structures in two configurations: top-down (root-to-leaf) and bottom-up (leaf-to-root).
+Figure 4 shows inefficiency across layers under both directions.
+In the top-down positioning, the energy dataset remains resilient up to layer 7 with 20% adversaries.
+Collapse occurs at layer 10 when the adversary ratio exceeds 50% and $\beta>0.9$.
+In voting, resilience holds through layer 4, with vulnerability at layer 5 and collapse at layers 6–7 under high severity ($\beta>0.9$).
+In the privacy datasets, both signals remain resilient in the top four layers, with vulnerability emerging at layer 4 for $\beta>0.7$.
+Collapse appears at layer 6 in the high privacy-preserving signal with over 60% adversaries, and slightly earlier in the low privacy-preserving signal with 50% adversaries.
+In the bottom-up direction, the energy dataset shows a narrower vulnerability and collapse region.
+Vulnerability begins at layer 9 with 50% adversaries, and collapse follows at layer 10 under high severity ($\beta>0.9$).
+Voting results are similar across both directions, with transitions driven more by adversary ratio than structural depth.
+Vulnerability emerges at layer 6 with 20% adversaries, and collapse follows at layer 5 with 60%.
+In privacy, vulnerability appears at layer 6 in both signals; collapse occurs at layer 6 (high privacy-preserving) and layer 5 (low privacy-preserving).
+(a) Inefficiency cost
+(b) Discomfort cost
+Figure 5: Normalized inefficiency and discomfort costs on synthetic Gaussian data across varying numbers of agents, plan options per agent, and adversary ratios.
+V-D Scalability Analysis with Synthetic Data
+To benchmark scalability and validate trends from real datasets, we use a synthetic Gaussian dataset to systematically vary agent population (10–100) and plan count (2–10) under different adversarial ratios.
+This controlled setup reveals optimization performance across configurations under adversarial pressure [^11] Results exclude $\beta=1$ to avoid saturation effects and highlight subtler trends..
+Figure 5(a) shows normalized inefficiency costs across agent counts, adversary ratios, and plan numbers.
+Inefficiency increases notably beyond 60% adversarial presence and is most pronounced in systems with smaller agent populations.
+Configurations with fewer plans exhibit high inefficiency even under moderate adversarial influence, indicating lower tolerance.
+At 100% adversaries, inefficiency peaks in 10-agent systems.
+In contrast, larger agent populations and broader plan spaces significantly enhance resilience, suppressing inefficiency across most adversarial scales.
+Figure 5(b) presents normalized discomfort costs under the same conditions.
+Discomfort increases with the number of plans.
+Systems with two plans show a gradual decline in discomfort as adversary ratios increase, while those with more plans experience sharper reductions.
+Discomfort variation expands with plan complexity, indicating greater sensitivity.
+The number of agents has minimal impact, though the 10-agent setting occasionally shows marginally lower discomfort than larger systems.
+V-E Summary of Findings and Discussion
+The key findings of this work can be summarized as follows:
+  1. 1.
+    The interplay of adversarial scale and severity determines the resilience, vulnerability and collapse of distributed multi-objective optimization, which is strongly influenced by the optimization scenario.
+  2. 2.
+    Distributed multi-objective optimization can predominantly remain resilient, even for high adversarial scales or severity.
+  3. 3.
+    Adversarial attacks trigger high comfort losses by legitimate agents as collective compromises that reduce the likelihood of collapse for higher vulnerability and resilience.
+  4. 4.
+    Comfort compromises of legitimate agents for preserving system efficiency under adversarial attacks are predominantly required for high adversarial severity.
+  5. 5.
+    Pareto optimal points for adversarial severity levels and adversarial scales are mainly in the resilience trajectory.
+However, Pareto optimal points for high adversarial scales can expand to vulnerability and collapse trajectories.
+  6. 6.
+    High adversarial scales reduce the comfort compromises required by the legitimate agents in the Pareto optimal points for adversarial severity levels that can be tolerated.
+  7. 7.
+    High adversarial severity levels reduce the system efficiency in the Pareto optimal points for adversarial scales that can be tolerated.
+  8. 8.
+    Lower hierarchical levels with higher scales of agents within hierarchical structures of distributed multi-objective optimization are more vulnerable to adversarial attacks than top levels with lower scales of agents.
+  9. 9.
+    A top-down positioning of adversary agents within hierarchical structures of distributed multi-objective optimization is more impactful on system performance: higher vulnerability, likelihood of collapse, and lower resilience.
+  10. 10.
+    Systems with a small number of agents and low plan diversity are more susceptible to inefficiency increases under adversarial pressure, even at moderate adversary ratios.
+  11. 11.
+    Broader plan spaces and larger agent populations enhance system resilience in distributed multi-objective optimization, significantly suppressing inefficiency across adversarial configurations.
+The findings of this paper can be used to develop and enhance corrective self-healing strategies [65] that are cost-effective in practice.
+They can also be used to design incentive mechanisms [66] that ensure agents comply to certain standards of safety in critical infrastructures such as smart grids.
+For instance, one challenge of fault-correction mechanism is the timely detection of adversary agents to mitigate their impact [65].
+Apparently, redundancy mechanisms and rollback operations orchestrated by monitoring mechanisms are resource-intensive [65, 67].
+They require frequent checks that involve computations and exchange of messages and they usually rely on static thresholds or even manual operations [67].
+This is where the insights of this work can find applicability: these mechanisms can adapt based on the status of the system, for instance, whether it is in the resilience, vulnerability and collapse state.
+Knowing apriori that an optimization process can tolerate certain adversarial scales and severity can simplify and reduce the costs of applying prevention and mitigation measures.
+It can also provide new insights for security policies, for instance, prioritizing the protection of agents at the top of the hierarchical optimization structure with stronger security safeguards and resources allocated for this purpose [68].
+While creating these strategies does not fall into the scope of this paper, it is part of the future work to pursue.
 VI Conclusion and Future work
-This study provides a comprehensive analysis of resilience, vulnerability, and collapse dynamics in multi-agent distributed optimization systems under adversarial conditions.
-By systematically examining the impact of the scale of adversaries, adversarial severity, and hierarchical structure, it identifies critical thresholds where system performance transitions from resilience to collapse.
-These insights offer practical strategies to enhance system design and resilience against adversarial influences.
-The results highlight critical thresholds where systems transition from resilience to collapse, emphasizing the interplay of structural vulnerabilities and adversarial factors.
-These findings provide a foundation for designing tailored resilience strategies across diverse operational contexts.
-A significant contribution of this work is the introduction of a benchmark dataset and software artifact for I-EPOS collective learning.
-These resources enable researchers to further investigate adversarial dynamics and develop advanced self-healing strategies.
-Integrating real-time detection and mitigation mechanisms will be critical for dynamically restoring system performance and resilience in the presence of adversarial behaviors.
-Future work will focus on embedding real-time adaptive mechanisms within practical systems to monitor and counteract adversarial influences dynamically.
-Exploring complex network topologies, diverse adversarial scenarios, and broader application domains will deepen understanding and support the development of robust, fault-tolerant distributed optimization systems capable of maintaining efficiency and stability under challenging conditions.
+This study provides a comprehensive analysis of resilience, vulnerability, and collapse dynamics in multi-agent distributed optimization under adversarial conditions.
+By systematically examining adversary scale, severity, and network structure, we identify critical thresholds where systems transition from stability to failure.
+These findings offer actionable guidance for designing and enhancing the performance of recovery and healing strategies.
+A key contribution of this work is the release of a large-scale benchmark dataset, generated from over 112 million experiments using the proposed adversarial model.
+This dataset supports systematic evaluation of adversarial impacts and facilitates reproducible research across domains of distributed optimization.
+Although the adversarial model is designed to be general-purpose, the evaluation can be extended to other algorithms in future work.
+Additionally, the current experiments model adversarial behavior with static severity levels, which may not fully capture dynamic or strategic adversary actions.
+Furthermore, resilience has been analyzed under hierarchical network structures, leaving the behavior under alternative topologies an open area for exploration.
+Future work will focus on embedding adaptive monitoring and mitigation mechanisms into real-time distributed systems.
+Investigating more complex network structures, dynamic adversarial strategies, and diverse application domains will further advance the development of robust, fault-tolerant optimization systems capable of maintaining performance under adversarial conditions.
 Acknowledgments
@@
 The authors would like to thank Thomas Wellings and Chuhao Qin for their feedback on the paper, Abhinav Sharma for technical support in experimentation and Srijoni Majumdar for support with the voting dataset.
+Appendix A: Cost Function Definitions
+This appendix presents the mathematical definitions of the cost functions used in the optimization experiments.
+The functions are inherited from the original I-EPOS framework [53] and are instantiated to reflect the characteristics of each application domain.
+Variance
+The variance cost measures the dispersion of the aggregated global response $g\in\mathbb{R}^{d}$ and is used in the energy and Synthetic Gaussian datasets. It is computed as:
+$f_{\text{var}}=\frac{1}{d}\sum_{j=1}^{d}\left(g_{j}-\bar{g}\right)^{2}$
+where:
+  • $g_{j}$ is the aggregated global response at dimension $j$,
+  • $\bar{g}$ is the mean of the global response across all dimensions,
+  • $d$ is the dimensionality of the response.
+Residual Sum of Squares (RSS)
+The RSS cost quantifies the squared difference between the scaled global response $g\in\mathbb{R}^{d}$ and a predefined system-wide target signal $T\in\mathbb{R}^{d}$.
+It is used in the voting and privacy datasets and is defined as:
+$f_{\text{RSS}}=\left(s(g)-s(T)\right)^{T}\left(s(g)-s(T)\right)$
+where $s(\cdot)$ denotes the scaling function applied to both vectors to improve shape alignment.
+Appendix B: Pareto Optimality Visualizations
+This appendix presents detailed visualizations of Pareto fronts for 10 selected adversarial severity levels ($\beta$) and 10 adversary population scales across the energy, voting, and privacy datasets.
+Each subfigure illustrates the trade-off between inefficiency cost and the discomfort cost of legitimate agents.
+Red lines indicate the non-dominated Pareto fronts, while red boxes mark the knee points, identified using the Minimum Manhattan Distance (MMD) method.
+These visualizations complement the analysis in Section V, offering deeper insights into system behavior under varying adversarial conditions.
+Appendix C: Pareto Optimality of Total Agents
+Figure 10 shows the trade‐off between system inefficiency and total‐agent discomfort, with Pareto knee points identified for both varying severity (a) and adversary scale (b).
+(a) Energy dataset — Pareto fronts and knee points across adversary population scales (10% to 100%).
+(b) Energy dataset — Pareto fronts and knee points across adversarial severity levels ($\beta=0.1$ to 1.0).
+Figure 6: Pareto optimality analysis for the Energy dataset.
+(a) Privacy dataset (high signal) — Pareto fronts and knee points across adversary population scales (10% to 100%).
+(b) Privacy dataset (high signal) — Pareto fronts and knee points across adversarial severity levels ($\beta=0.1$ to 1.0).
+Figure 7: Pareto optimality analysis for the high privacy-preserving signal in the privacy dataset.
+(a) Privacy dataset (low signal) — Pareto fronts and knee points across adversary population scales (10% to 100%).
+(b) Privacy dataset (low signal) — Pareto fronts and knee points across adversarial severity levels ($\beta=0.1$ to 1.0).
+Figure 8: Pareto optimality analysis for the low privacy-preserving signal in the privacy dataset.
+(a) Voting dataset — Pareto fronts and knee points across adversary population scales (10% to 100%).
+(b) Voting dataset — Pareto fronts and knee points across adversarial severity levels ($\beta=0.1$ to 1.0).
+Figure 9: Pareto optimality analysis for the voting dataset.
+(a) Pareto front and knee points across adversarial severity.
+(b) Pareto front and knee points across scales of adversaries.
+Figure 10: The Pareto Optimality of The energy voting, and privacy datasets